### v1

# RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes

Abstract
Accurate detection of changes in roads and bridges, such as construction, renovation, and demolition, is essential for urban planning and traffic management. However, existing methods often struggle to extract fine-grained semantic change information due to the lack of high-quality annotated datasets in traffic scenarios. To address this, we introduce the Road and Bridge Semantic Change Detection (RB-SCD) dataset, a comprehensive benchmark comprising 260 pairs of high-resolution remote sensing images from diverse cities and countries. RB-SCD captures 11 types of semantic changes across varied road and bridge structures, enabling detailed structural and functional analysis. Building on this dataset, we propose a novel framework, Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain. MFDCD includes a Dynamic Frequency Coupler (DFC) that fuses hierarchical visual features with wavelet-based frequency components, and a Textual Frequency Filter (TFF) that transforms CLIP-derived textual features into the frequency domain and applies graph-based filtering. Experimental results on RB-SCD and three public benchmarks demonstrate the effectiveness of our approach.
keywords:
[…]
1 Introduction
In the field of remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes (Pekel et al., 2016; Song et al., 2018). With the acceleration of urbanization and the continuous development of transportation infrastructure, CD in traffic scenes is crucial for traffic planning and facility management. In traffic planning, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades (Lin and Habib, 2022; K S et al., 2024; Wang et al., 2025). For instance, the emergence of new overpasses, flyovers, or road expansions needs to be timely integrated into transportation maps to ensure up-to-date routing and navigation services. Moreover, detecting illegal road constructions or temporary blockages is also critical for enforcing urban regulations and minimizing congestion. From the perspective of traffic facility management, road and bridge CD enables timely identification of structural deterioration, unauthorized demolitions, or unreported construction activities. This semantic-level understanding is crucial for preventive maintenance, disaster response (e.g., after floods or earthquakes), and resource allocation. Additionally, as smart cities and autonomous transportation systems increasingly rely on accurate and up-to-date digital maps, the need for automated and fine-grained CD becomes even more pressing. Roads and bridges are not only physical assets but also key elements in the digital infrastructure of intelligent transportation systems.
However, as shown in Table  1, traditional binary change detection (BCD) datasets (Ji et al., 2018; Chen and Shi, 2020; Shi et al., 2021), etc. primarily focus on building CD and only provide simple information about whether a change has occurred, represented by 0 and 1 for no change and change, respectively. On the other hand, existing semantic change detection (SCD) datasets (Caye Daudt et al., 2019; Yuan et al., 2022; Yang et al., 2022a), etc. mainly focus on changes in land cover, such as building changes and vegetation changes. Changes related to roads and bridges in traffic scenes are either limited in number or completely absent. To address this issue, as annotation examples shown in Figure 1, we propose a new dataset specifically designed for SCD of roads and bridges in traffic scenes: Road and Bridge Semantic Change Detection (RB-SCD) dataset. RB-SCD provides semantic annotations for 11 distinct categories of road and bridge changes, effectively addressing insufficient granularity of information and lack of specificity in detecting road and bridge changes in traffic scenes.
In addition to the RB-SCD dataset, we have also developed a new framework to further facilitate CD performance, the Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain. Considering CD methods, traditional approaches often rely on convolutional neural network architectures (Caye Daudt et al., 2018; Fang et al., 2022; Feng et al., 2022), which are insufficient to fully exploit the rich spatiotemporal information in images. Some studies (Chen et al., 2022; Bandara and Patel, 2022; Feng et al., 2023) have incorporated transformers into the network for global modeling. However, these methods still face challenges in accurately capturing and effectively distinguishing complex textures and structural changes. Research has also shown that relying solely on single-modal data has inherent limitations. Recently, some methods  (Yang et al., 2022b; Zhang et al., 2023; Dong et al., 2024) have introduced multimodal features to overcome the shortcomings of single modalities, such as combining visual and textual information or integrating lidar and hyperspectral data. However, these approaches often fail to fully exploit the characteristics of multimodal features, focusing only on modeling the commonalities of multimodal data while neglecting the distributional differences among different modalities.
Table 1: The public datasets for urban change detection with optical remote sensing imagery.
Dataset Resolution (m) Image pairs Image size (pixels) Change type Objects of interest Classes
LMR (Low-to-Medium Resolution)
OSCD  (Daudt et al., 2018) 10 24 $600\times 600$ Binary Land cover
S2MLTCP  (Leenstra et al., 2021) 10 1520 $600\times 600$ Binary Land cover
ZY3  (Zhang and Shi, 2020) 5.8 1 $1154\times 740$ Binary Land cover
Landsat-SCD  (Yuan et al., 2022) 30 8468 $416\times 416$ Semantic Land cover 4
HR (High-Resolution)
WV3  (Zhang and Shi, 2020) 2 2 $1431\times 1431$ Binary Land cover
QB  (Zhang and Shi, 2020) 2.4 1 $1154\times 740$ Binary Land cover
VHR (Very High-Resolution)
AICD  (Bourdis et al., 2011) 0.5 1000 $800\times 600$ Binary Land cover
LEVIR-CD  (Chen and Shi, 2020) 0.5 637 $1024\times 1024$ Binary Buildings
WHU  (Ji et al., 2018) 0.2 1 $32207\times 15354$ Binary Buildings
CDD  (Lebedev et al., 2018) 0.03 - 1 16000 $256\times 256$ Binary Buildings
SYSU  (Shi et al., 2021) 0.5 16000 $256\times 256$ Binary Buildings
BDD  (Adriano et al., 2021) 1.0 1147 $250\times 250$ Semantic Buildings 1
HRSCD  (Caye Daudt et al., 2019) 0.5 291 $10000\times 10000$ Semantic Land cover 5
SECOND  (Yang et al., 2022a) 0.5 - 3 4662 $512\times 512$ Semantic Land cover 6
Hi-UCD  (Tian et al., 2020) 0.1 1293 $1024\times 1024$ Semantic Land cover 9
RB-SCD (Ours) 0.59 260 $270\times 534$ $7215\times 4366$ Semantic Roads and bridges 8
Figure 1: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represents pre- and post-change images, respectively.
Figure 2: Visualization of visual and textual frequency spectrum.
In this work, as shown in Figure 2, we observe that the decomposition and utilization of features in the frequency domain remain underexplored, particularly the fine-grained change information embedded in the high-frequency visual features, the large-scale distribution information in the low-frequency features, and the frequency-domain representation of textual features. The potential patterns of multimodal information in the frequency domain have not been adequately studied. Therefore, we make the first attempt to develop a new architecture that maps image and textual features into different types of frequency domains based on their characteristics, enabling multimodal frequency-domain modeling for CD.
Our MFDCD consists of two main components: Dynamic Frequency Coupler (DFC) and the Textual Frequency Filter (TFF). DFC decomposes bi-temporal features into high- and low-frequency components using discrete wavelet transform (DWT), and introduces an Adaptive Sparse Frequency Fusion (ASFF) block for the combination of high-frequency information and low-level features. During the early stages of training, ASFF enhances the sparsity of high-frequency features, focusing on significant change areas. In the later stages, the sparsity constraint is reduced to enhance the expressive power of the model, ultimately increasing the attention to key change areas while suppressing high-frequency noise interference. In this way, edge and texture information in low-level features can not only be gradually captured through convolution operations but also dynamically weighted based on the localized changes in refined high-frequency information. For low-frequency features, Bidirectional Temporal Frequency Fusion (BTFF) block is introduced. Since CD involves not only capturing local changes but also understanding the relative changes between different time points, simply combining the high-level features at a single time point with low-frequency features could cause the model to overly rely on the local structural information at that moment, neglecting the cross-temporal changes. The BTFF block addresses this by combining high-level features with low-frequency features across time points, enabling the model to understand the trend of changes and global structural changes over time.
TFF introduces a pre-trained CLIP (Radford et al., 2021) model to obtain textual information, which is further expanded to get text descriptors. Text features contain semantic information but lack spatial features, and discrete Fourier transform (DFT) can convert them into frequency domain information, making it easier to align with the frequency information of image features. By constructing graph filters to aggregate key information, the processed textual features are deeply integrated with visual features to provide multimodal support for SCD.
Our contributions are as follows: {adjustwidth}1.5em0pt•  A comprehensive benchmark for SCD in traffic scenarios. We construct a new dataset named RB-SCD, which focuses on SCD of roads and bridges. It contains 260 large-scale high-resolution RS image pairs, covering over ten cities across multiple countries and annotating 11 typical semantic change types. With its wide geographic diversity and complex traffic infrastructures, RB-SCD supports fine-grained structural and functional change analysis for transportation environments. {adjustwidth}1.5em0pt•  A novel multimodal frequency-driven CD framework. We propose Multimodal Frequency-Driven Change Detector (MFDCD), the first framework to introduce frequency-domain multimodal feature integration for CD. It consists of two main modules: (1) Dynamic Frequency Coupler (DFC), which fuses hierarchical visual features with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception; (2) Textual Frequency Filter (TFF), which transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative multimodal representation learning. {adjustwidth}1.5em0pt•  Comprehensive experiments on multiple benchmarks. Extensive experiments on the proposed RB-SCD and three public CD datasets demonstrate that MFDCD achieves superior performance and generalization ability in identifying complex semantic changes in traffic scenes.
2 Related Work
2.1 CD Datasets and Methods
Existing datasets, as shown in Table 1, such as  (Ji et al., 2018; Lebedev et al., 2018; Chen and Shi, 2020; Shi et al., 2021), are primarily designed for BCD tasks and cannot provide specific semantic change information. For SCD, datasets like  (Caye Daudt et al., 2019; Yuan et al., 2022; Yang et al., 2022a) have established some benchmarks. However, these datasets mainly focus on changes in land cover, with little attention paid to the changes in roads and bridges in traffic scenes. In contrast, our RB-SCD dataset is a high-resolution, fine-grained SCD dataset specifically curated and annotated for road and bridge changes in traffic scenes. It captures a wide range of semantic change types and includes diverse geographical contexts, enabling precise and meaningful evaluation of CD algorithms in real-world urban transportation scenarios. By addressing both the lack of semantic detail and the absence of domain specificity in existing datasets, RB-SCD provides a valuable and targeted benchmark for infrastructure monitoring, intelligent transportation analysis, and traffic scene understanding.
Figure 3: (a) Statistical distribution of pixels for semantic change categories. (b) Train set. (c) Test set.
Table 2: Specific subcategories in RB-SCD dataset.
Categories Description
Road street, highway, dirt road, walkway, railway, etc
Bridge
beam bridge, arch bridge, cable-stayed bridge,
suspension bridge, etc.
Bare land
wasteland, sand patches area, unpaved open
grounds, road or railway construction zones, etc.
Grass
urban lawn, green belt, wetland grass, riparian
grass zone, meadow etc.
Water
ocean, lake, river, reservoir, pond, etc.
Farmland agricultural land (with and without crops)
Vegetation forest, shrub, urban vegetation buffer, etc.
Building
residential building, commercial building,
industrial building, under-construction building
(with clear structure), etc.
In recent years, deep learning has significantly improved the accuracy of CD through its ability to extract relevant features from raw data. Changemask (Zheng et al., 2022) decouples SCD into temporal semantic segmentation and BCD, and combines these two tasks into a unified encoder-transformer-decoder framework. MDENet (Liu et al., 2024) uses MDCM to capture object features in frequency and spatial domains across times, and MFGatt to reduce style impact, focusing on object location. JFRNet (Chang et al., 2024) introduces a three-branch network that takes joint bi-temporal features as unified input, enabling each branch to perceive the semantic relevance of the two time points without any additional interaction operations. DecoderFusion (Li et al., 2024) enhances CD decoding features by leveraging the feature differences between two semantic segmentation decoding branches. CdSC (Wang et al., 2024) introduces a cross-differential semantic consistency network that explores the differences in bi-temporal features while maintaining their semantic consistency. HGINet (Long et al., 2024) builds a multi-level perception aggregation network with a pyramid architecture to extract features that distinguish different categories across multiple levels. However, these methods only consider single-modal data, i.e., images, while neglecting the rich semantic information embedded in multimodal data, thus encountering a performance bottleneck.
2.2 Multimodal Learning
Recent studies have shown that multimodal information is beneficial for visual image analysis. LAVT (Yang et al., 2022b) and VLT (Ding et al., 2023) integrate textual features into visual features through language encoders. Zhang et al. (Zhang et al., 2023) propose a convolutional neural network method with an attention mechanism to enhance feature extraction for lidar data. Their approach leverages the elevation information from lidar data and the spatial-spectral information of hyperspectral data for multimodal feature fusion. MedCLIP (Koleilat et al., 2024) uses text prompts to generate segmentation for clinical scans, adopting a novel decoupled hard-negative noise contrastive estimation loss to fine-tune the BiomedCLIP model. ChangeCLIP (Dong et al., 2024) designs a novel CD framework that utilizes semantic information from image-text pairs and introduce an innovative differential feature compensation module to capture fine-grained semantic changes between them.
2.3 Frequency Transform in Deep Learning
In image processing, the combination of CNN and frequency domain information has been widely applied in tasks such as image restoration, super-resolution, and semantic segmentation. W-Unet (Zhao et al., 2021) replaces down-sampling operations with DWT to reduce the impact of image noise while avoiding information loss. FFT-ReLU  (Mao et al., 2023) performs image deblurring tasks in the frequency domain based on fast Fourier transform. ARFFT (Zhu et al., 2023) proposes a spatial-frequency fusion block to enhance the representation capability of Transformers and extend the receptive field across the entire image. XNet (Zhou et al., 2023) utilizes wavelet transforms to transfer visual information to the frequency domain, leveraging high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks. Given the characteristics of RS images, such as high resolution and susceptibility to spectral variations caused by seasonal changes, using only frequency domain features often results in the loss of critical spatial information. Therefore, in this paper, we propose a novel architecture that models multimodal features in the frequency domain, effectively balancing the advantages of visual features, textual features, and frequency domain features, achieving promising results.
Figure 4: Examples of dataset distribution spanning multiple representative geographical areas.
Figure 5: Examples of road and bridge types, where the first two columns are roads, and the last three columns are bridges.
3 RB-SCD Dataset
3.1 Data Annotation and Analysis
Rich Semantic Annotations: As shown in Figure 1 and Figure 3, unlike traditional BCD datasets, we have defined eight major categories (with specific subcategories detailed in Table 2 ), including road, bridge, bare land, building, farmland, water, grass and vegetation. The RB-SCD dataset provides detailed semantic change annotations for 11 distinct types of semantic changes and a background category across these eight categories.
Figure 6: Overview of our proposed Multimodal Frequency-Driven Change Detector (MFDCD).
Changes such as road widening, new road construction, and bridge construction or modification are reflected through the corresponding semantic change types, supporting in-depth exploration of the intrinsic mechanisms of traffic scenario evolution. It is worth noting that some changes, such as bridges directly turning into roads, are rarely observed in reality. Additionally, other types of changes, such as those between bridges and some other categories, occur in only a few samples, and we set their pixel number to 0. All annotations are carried out by a professional team in RS and transportation, and underwent multiple rounds of rigorous review and quality control to ensure annotation accuracy and semantic consistency.
Diverse Geographic Coverage: As shown in Figure 4, following the principles in (Long et al., 2021), we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set. The dataset integrates satellite images obtained primarily from Google Earth, covering various regions worldwide, including Chinese cities such as Hefei, Xi’an, Chongqing and Guangzhou, as well as representative areas from Japan, Singapore, Nigeria, Italy, America, and Indonesia, etc. This extensive geographic coverage ensures that the dataset encompasses traffic scenes under varying climatic conditions, geological structures, and urban planning contexts, thereby guaranteeing its diversity and representativeness.
Multiple Road and Bridge Types: As shown in Figure 5, the dataset encompasses various types of roads and bridges from different geographical regions and urban environments. The road categories include concrete and asphalt roads used for urban main roads, secondary roads, residential streets, and highways, comprehensively reflecting the common transportation routes found in cities. Additionally, the dataset contains elevated roads designed to alleviate congestion and improve traffic efficiency, which are typically found in densely populated or high-traffic urban areas. In terms of bridge structures, the dataset includes large cable-stayed bridges that span wide rivers or deep valleys, as well as girder bridges and pedestrian bridges used for short- to medium-span connections and pedestrian access. The dataset also includes other types of specialized roads and bridges, demonstrating a variety of engineering styles and applications. This richness in road and bridge types enables the dataset to represent complex traffic environments and provides robust support for training CD models across multiple scales and diverse scenarios.
3.2 Label Imbalance Analysis and Challenges
As shown in Figure  3, our RB-SCD dataset exhibits a certain degree of sample imbalance, primarily manifested in the unequal distribution of different road and bridge types as well as change categories. Due to the diversity in data collection times and geographical regions, some categories of roads, bridges, and change types have fewer samples. In terms of semantic changes, there is also an imbalance in the types of changes in the dataset, especially for fine-grained CD tasks. For some types of changes (e.g., road widening, bridge construction), there are more samples, while for other types of changes (e.g., road repair, demolition), there are fewer samples. This imbalance may lead to the model overfitting to the more common change types, while neglecting the rarer ones. Additionally, the imbalance between change and no-change samples is another issue, particularly for CD tasks. The no-change samples are relatively more abundant, which may lead to a decrease in the model’s sensitivity when processing change samples. To improve the accuracy and robustness of the model, we will further collect more samples of different types of changes in future work and employ techniques such as data augmentation and sample weighting to alleviate the model’s difficulty in detecting rare changes.
4 Methodology
4.1 Overview
As shown in Figure 6, our MFDCD mainly includes two components: (1) DFC module effectively couples high-level and low-level visual features with high-frequency and low-frequency spectral features. (2) TFF module enhances the text information obtained through zero-shot inference, explores the representation of text features in the frequency domain, and promotes the interaction and fusion of multimodal data. The details are as follows.
4.2 DFC Module
Given the bi-temporal images $T_{1}$ and $T_{2}$, ResNet50 (He et al., 2016) is used as the backbone network to extract the corresponding multi-scale features $F_{i}^{t}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2)\right\}$. The features from the first two layers of the network are considered as low-level features, while the features from the last two layers are considered as high-level features. The low-frequency and high-frequency components of these features are then obtained through wavelet transform, which allows for the separation of coarse and fine-grained information in both spatial and frequency domains.
Adaptive Sparse Frequency Fusion: Low-level features are rich in detailed information, while high-frequency features contain information about edges, details, and abrupt changes, which naturally correspond to low-level features. However, high-frequency features are sensitive to noise and may amplify the influence of irrelevant regions. As shown in the upper-right corner of Figure  6, the ASFF block applies a novel sparsity constraint to the high-frequency features, which is dynamically adjusted during training iterations. By combining the sparsity-constrained high-frequency features with low-level features, both the image’s detailed information (such as edges and textures) and local structural information are preserved. This combination makes the network more responsive to CD, especially in areas with detailed changes. The process is as follows:
$LL_{i}^{t},HL_{i}^{t},LH_{i}^{t},HH_{i}^{t}=\mathcal{W}(F_{i}^{t}),$ (1)
where $t\in\left\{1,2\right\}$, $i\in\left\{1,2,3,4\right\}$, $\mathcal{W}(\cdot)$ represents the discrete wavelet transform. $LL_{i}^{t}$ is the low-frequency band, $HL_{i}^{t}$, $LH_{i}^{t}$, and $HH_{i}^{t}$ signifying the high-frequency bands.
$H_{i}^{t}=Conv(Cat(HL_{i}^{t},LH_{i}^{t},HH_{i}^{t})),$ (2)
$\widetilde{H_{i}^{t}}=\mathcal{S}(H_{i}^{t},\lambda(p),g)=\begin{Bmatrix}H_{i}^{t},&if\left|H_{i}^{t}\right|>g\cdot\lambda(p)\\ 0,&otherwise\end{Bmatrix},$ (3)
$\lambda(p)=\lambda_{0}\cdot e^{-\gamma p},$ (4)
$P_{i}^{t}=Conv(Cat(Up(\widetilde{H_{i}^{t}}),F_{i}^{t})),$ (5)
where $t\in\left\{1,2\right\}$, $i\in\left\{1,2\right\}$. $\mathcal{S(\cdot)}$ denotes the sparsity constraint function, which is used to apply the sparsity operation, $\lambda(p)$ is the dynamic sparsity strength, $\lambda_{0}$ is the initial sparsity strength, $\gamma$ is the sparsity decay rate, $p$ is the current training iteration, and $g$ is a constant controlling the sparsity strength. $H_{i}^{t}$ is the original fused high-frequency feature and $\widetilde{H_{i}^{t}}$ is the high-frequency feature after applying the sparsity constraint. $P_{i}^{t}$ is the enhanced feature that combines high-frequency and low-level information. $Conv(\cdot)$ and $Up(\cdot)$ are convolution and bilinear upsampling operation, respectively.
Bidirectional Temporal Frequency Fusion: As shown in the lower-right corner of Figure 6. Take the low-frequency features of time $T_{2}$ and the high-level features of time $T_{1}$ as an example, high-level features at time $T_{1}$ typically contain the semantic hierarchy of the image, enabling the expression of global structural changes in the image, such as road widening and bridge construction. Low-frequency features at time $T_{2}$ mainly focus on smooth regions or large-scale change information in the image. By combining high-level features from $T_{1}$ with low-frequency features from $T_{2}$, the network can capture global structural changes across time points while maintaining sensitivity to large-scale, smooth region changes. This models the bidirectional relationship of changes: local changes at $T_{1}$ may lead to semantic-level changes at $T_{2}$, while detailed changes at $T_{2}$ may arise from regional distribution adjustments at $T_{1}$. The details of BTFF block are as follows:
$P_{i}^{1}=Conv(Cat[Up(LL_{i}^{2}),F_{i}^{1}]),$ (6)
$P_{i}^{2}=Conv(Cat[Up(LL_{i}^{1}),F_{i}^{2}]),$ (7)
where $i\in\left\{3,4\right\}$, $P_{i}^{t}$ is the enhanced feature that combines low-frequency information and high-level information.
4.3 TFF Module
As shown in Figure 6, we leverage the powerful zero-shot inference capability of the CLIP model to obtain textual information from RS images for multimodal modeling. For the categories inferred by the CLIP model, we select the top five categories with the highest probability values. For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ," where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself. $E_{T}$ is a pre-trained CLIP text encoder that encodes the text embedding. Considering that discrete Fourier transform can help to reveal high-order semantic patterns in text, especially by identifying more complex semantic information through the descriptors used at different temporal points ($T_{1}$, $T_{2}$). Through Fourier transform, periodic or repetitive structures in the text can be extracted, which helps identify commonalities and differences in changes. We first use Fourier transform to convert the text embeddings from the text domain to the frequency domain:
$\mathbf{T}_{freq}=\mathcal{F}(E_{T}(T_{d})),$ (8)
where $\mathcal{F}(\cdot)$ is the discrete Fourier transform, $\mathbf{T}_{freq}\in\mathbb{C}^{B\times N\times C}$, $B$ represents the batch size, $N$ represents the number of frequency components, $C$ represents the dimension of each frequency component. We introduce a graph filter bank $K=[k_{1},k_{2},...k_{k}]$ for the textual features, which propagate information between the nodes of the graph to filter and refine the frequency information, thereby obtaining salient features related to the changes. For the frequency feature $\mathbf{T}_{freq}$, we split it into real and imaginary parts and concatenate them to construct node features. The feature matrix $\mathbf{X}\in\mathbb{R}^{B\times N\times 2C}$ is the collection of features for all nodes, which allows us to convert frequency domain information into node features for the graph. Assuming that each frequency component is connected to every other frequency component (fully connected graph), the edge set $\mathbf{E}\in\mathbb{Z}^{2\times\frac{N(N-1)}{2}}$ of the graph can be constructed by combining each pair of nodes:
$\mathbf{X}=Cat[\mathfrak{R}(\mathbf{T}_{freq}),\mathfrak{I}(\mathbf{T}_{freq})],$ (9)
$\mathbf{E}=\left\{(i,j)\mid 1\leq i\neq j\leq N\right\},$ (10)
$\mathbf{E}^{(k+1)}=\sigma(\hat{\mathbf{A}}\mathbf{E}^{(k)}\mathbf{W}_{k}),$ (11)
$\mathbf{E}^{f}=\sum_{k=1}^{k}\mathbf{E}^{(k)},$ (12)
where $\mathfrak{R}(\mathbf{T}_{freq})$ and $\mathfrak{I}(\mathbf{T}_{freq})$ represent the real and imaginary parts of the frequency domain features, respectively. $\mathbf{E}^{(k)}$ represents the node features after the $k$-th filter, $\hat{\mathbf{A}}$ is the normalized adjacency matrix, which represents the relationships between nodes, $\mathbf{W}_{k}$ is the weight matrix of the graph convolution filter at the $k$-th layer, and $\sigma$ is the activation function, $\mathbf{E}^{f}$ is the weighted sum of the outputs from multiple filters. We use the text-image encoder $E_{VT}$ of the CLIP model to fuse $\mathbf{E}^{f}$ and the image features. After subtracting the results from the bi-temporal phases, the output is fed into the decoder (Chen et al., 2024) to obtain our CD results.
4.4 Frequency Transform Function Analysis
Why use discrete wavelet transform for image features and discrete Fourier transform for textual features? The advantage of wavelet transform lies in its ability to analyze images at different scales and locations, making it particularly suitable for multi-level, multi-scale image analysis tasks. As shown in Figure  2, in remote sensing image CD, wavelet transform helps to extract high-frequency features (e.g., detailed changes) and low-frequency features (e.g., overall structure), aiding in the capture of changes at different spatiotemporal scales in the image. Considering that text is typically discrete sequential data, its changes are often global or frequency-related. Therefore, on the other hand, the Fourier transform can effectively reveal periodic semantic patterns in the text, which contributes to identify patterns and trends in text changes.
5 Experiments and Results
5.1 Public CD Datasets
In addition to using our proposed dataset, as shown in Table 3, we have employed three public CD datasets to evaluate the robustness of our model. The LEVIR-CD (Chen and Shi, 2020) dataset consists of 637 pairs of RS images with a resolution of 1,024 × 1,024. It mainly focuses on building area changes, such as building growth. The WHU-CD (Ji et al., 2018) dataset comprises a pair of high-resolution RS images with a size of 32,507 × 15,354 pixels. It covers areas before and after an earthquake and primarily includes building renovations and additions. The SYSU-CD (Shi et al., 2021) dataset is a collection of 20,000 pairs of high-resolution aerial images, each with a size of 256 × 256 pixels. The images are captured in the Hong Kong area and include urban construction, suburban expansion, and vegetation changes, etc.
Table 3: Public CD datasets used in the experiments.
Datasets Train/ Val/ Test Image type Space resolution
LEVIR-CD 7120/ 1024/ 2048 Satellite images 0.5 m/pixel
SYSU-CD 12000/ 40000/ 4000 Aerial images 0.5 m/pixel
WHU-CD 5947/ 743/ 744 Aerial images 0.2 m/pixel
Table 4: Comparison of quantitative performance for semantic change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Class HGINet (Long et al., 2024) CdSC (Wang et al., 2024) DEFO (Li et al., 2024) ChangeCLIP (Dong et al., 2024) Ours
IoU F1 Rec Pre IoU F1 Rec Pre IoU F1 Rec Pre IoU F1 Rec Pre IoU F1 Rec Pre
background 97.76 98.87 99.14 98.60 98.22 99.1099.45 98.75 98.18 99.08 98.27 98.90 98.11 99.05 99.37 98.73 98.28 99.13 99.34 98.93
bare land ->road 45.24 62.30 61.10 63.54 48.42 65.25 63.40 67.21 50.93 67.49 66.0469.00 51.75 68.21 64.87 64.87 52.76 69.0869.4968.67
water ->bridge 80.20 89.01 85.60 92.71 85.33 92.09 93.32 90.88 83.10 90.77 91.38 90.17 85.77 92.34 88.22 96.87 86.83 92.9594.28 91.67
water->road 58.09 73.49 76.88 70.39 68.64 81.49 74.82 89.26 62.78 77.14 76.21 78.88 68.11 81.03 76.23 86.47 71.68 83.51 79.6387.78
road->bare land 33.83 50.55 40.13 68.29 38.33 55.42 44.58 73.22 38.14 55.22 48.02 64.94 40.57 57.7247.0674.63 39.48 56.61 43.88 79.73
road->vegetation 10.69 19.32 12.56 41.82 13.65 24.03 15.25 56.57 17.02 29.0922.47 41.25 16.37 28.13 19.35 51.49 20.31 33.7621.8374.39
vegetation->road 52.61 68.94 63.10 75.98 57.93 73.36 69.89 77.19 57.84 73.29 72.30 74.41 57.10 72.69 68.55 77.37 59.35 74.49 72.38 76.72
building->road 36.56 55.57 43.44 69.78 36.10 53.05 41.28 74.18 39.83 56.97 51.63 63.54 40.46 57.61 44.69 81.04 45.44 62.49 53.0675.98
farmland->road 63.36 73.55 77.80 77.34 63.80 77.90 78.78 86.61 65.38 79.07 75.87 82.55 65.02 78.80 81.30 76.45 68.40 81.24 79.21 83.38
grass->road 36.55 53.54 39.09 84.94 36.66 53.65 46.45 64.48 40.60 57.7548.26 71.89 36.55 53.54 39.09 84.94 42.49 59.6447.24 80.86
road->grass 22.44 36.65 32.38 42.21 15.53 26.89 17.41 58.97 18.41 31.10 22.75 49.14 21.76 35.7524.2168.26 21.65 35.59 22.86 80.30
bridge->water 46.51 63.49 49.92 87.19 79.36 82.60 74.25 93.06 83.40 90.9593.45 88.57 77.86 87.55 90.44 84.85 83.58 91.05 90.34 91.78
mIoU 48.65 52.75 54.63 54.95 57.52
Table 5: Performance of the RS foundation model LWGANet on RB-SCD dataset. All metrics are expressed as percentages (%).
Class LWGANet-L1 (Lu et al., 2025) LWGANet-L2 (Lu et al., 2025)
IoU F1 Pre Rec IoU F1 Pre Rec
background 98.3 99.14 98.84 99.44 98.27 99.13 98.8 99.45
bare land ->road 54.17 70.27 73.28 67.5 53.17 69.42 74.46 65.02
water ->bridge 87.29 93.21 92.52 93.92 85.85 92.39 90.95 93.87
water->road 69.53 82.03 82.51 81.55 63.42 77.62 83.84 72.25
road-> bare land 38.54 55.64 78.94 42.95 38.06 55.14 74.67 43.71
road->vegetation 15.54 26.89 55.44 17.75 13.7 24.1 68.9 14.6
vegetation->road 59.14 74.32 75.55 73.13 60.1 75.08 78.81 71.68
building->road 42.1 59.25 81.97 46.4 42.27 59.42 76.5 48.57
farmland->road 69.94 82.31 86.16 78.79 71.04 83.07 83.96 82.2
grass->road 28.98 44.93 81.75 30.98 40.78 57.93 81.93 44.81
road->grass 22.78 37.11 83.44 23.86 21.87 35.89 84.47 22.79
bridge->water 75.1 85.78 89.44 82.41 80.01 88.9 89.56 88.24
mIoU 55.12 55.71
Table 6: Comparison of quantitative performance for binary change detection on RB-SCD dataset. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Model IoU F1 Rec Pre
FC-diff (Caye Daudt et al., 2018) 57.81 73.26 69.63 77.30
FC-ef (Caye Daudt et al., 2018) 59.05 74.26 69.05 80.31
FC-cat (Caye Daudt et al., 2018) 60.34 75.26 71.99 78.85
SNUNet (Fang et al., 2022) 62.90 77.22 72.47 82.64
BIT (Chen et al., 2022) 62.21 76.71 70.97 83.69
ChangeFormer (Bandara and Patel, 2022) 61.88 76.45 69.69 84.66
ICIF (Feng et al., 2022) 61.80 76.39 71.93 81.43
TFI-GR (Li et al., 2022) 64.72 78.58 73.78 84.05
DMINet (Feng et al., 2023) 62.99 77.29 72.14 83.24
USSFC-Net (Lei et al., 2023) 62.55 76.96 75.07 78.95
ELGC-Net (Noman et al., 2024) 65.53 79.18 73.99 85.15
Ours 65.83 79.39 74.13 85.47
Table 7: Comparison of quantitative performance of various CD techniques on public datasets. All metrics are expressed as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance. Results marked with * are obtained from the original papers.
Model LEVIR-CD (Chen and Shi, 2020) SYSU-CD (Shi et al., 2021) WHU-CD (Ji et al., 2018)
IoU  F1  Rec  Pre IoU  F1  Rec  Pre IoU  F1  Rec  Pre
FC-diff (Caye Daudt et al., 2018) 75.18 85.83 84.59 87.12 42.05 59.20 46.72 80.80 69.48 81.99 78.90 89.14
FC-ef (Caye Daudt et al., 2018) 71.80 83.59 82.27 84.94 55.83 71.65 65.41 79.22 58.53 71.65 65.41 79.22
FC-cat (Caye Daudt et al., 2018) 76.98 86.99 85.64 88.38 60.68 75.53 71.38 80.21 72.24 83.87 84.22 83.53
SNUNet (Fang et al., 2022) 82.20 90.23 88.97 91.54 66.73 80.04 79.79 80.30 76.02 86.38 88.95 83.95
BIT (Chen et al., 2022) 82.06 90.15 88.98 91.34 62.50 76.93 72.50 81.93 80.40 89.13 89.69 88.58
ChangeFormer (Bandara and Patel, 2022) 81.93 90.07 87.38 92.92 60.46 75.36 71.06 80.22 83.88 91.23 90.76 91.71
ICIF (Feng et al., 2022) 79.86 88.80 86.85 90.84 61.38 76.07 73.31 91.94 78.27 87.81 84.06 91.96
TFI-GR (Li et al., 2022) 83.03 90.73 89.53 91.95 71.38 83.30 81.16 85.56 88.55 93.92 92.65 95.24
DMINet (Feng et al., 2023) 83.14 90.79 88.89 92.78 68.20 81.10 76.82 85.88 79.53 88.60 91.24 86.10
USSFC-Net (Lei et al., 2023) 82.54 90.44 91.82 89.09 65.75 79.34 81.52 77.27 83.92 91.26 93.20 89.40
M-Swin * (Pan et al., 2024) 83.57 91.05 90.05 92.07 -   -   -   - 86.09 92.94 91.37 94.56
ELGC-Net (Noman et al., 2024) 83.52 91.02 89.78 92.30 65.34 79.03 77.26 80.89 75.56 86.08 82.15 90.40
MDENet * (Liu et al., 2024) 83.63 91.0992.68 89.55 -   -   -   - 87.74 93.47 95.49 91.54
Ours 84.69 91.7192.31 91.11 72.63 84.15 85.60 82.74 90.43 94.97 96.44 93.55
5.2 Implementation Details
Our experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU. The AdamW (Kingma and Ba, 2014) is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01. Cross entropy loss is used as our loss function, and the maximum number of iterations is set to 40,000. For data augmentation, RandomFlip and PhotoMetricDistortion are applied. The constant $g$ in the ASFF block is set to 0.1, the decay rate $\gamma$ is set to 0.0001, and $\lambda_{0}$ is set to 1. To evaluate the performance of our experiments, we adopt four widely used metrics: recall (Rec), precision (Pre), F1-score (F1) and intersection over union (IoU), where IoU and F1 are the primary metrics, while the others are auxiliary metrics. The following are formulas for defining the evaluation metrics as mentioned above.
$Rec=TP/(TP+FN),$ (13)
$Pre=TP/(TP+FP),$ (14)
$F1=2/(Rec^{-1}+Pre^{-1}),$ (15)
$IoU=TP/(TP+FP+FN),$ (16)
where TP, TN, FP, and FN denote the counts of true positives, true negatives, false positives, and false negatives, respectively.
5.3 Quantitative Assessment Results
Results for Semantic Change Detection on RB-SCD Dataset: Table  4 presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics. As shown in the table, we have defined 11 types of semantic changes along with one background category, aiming to comprehensively cover the diverse real-world scenarios of road and bridge changes. These change types include both relatively straightforward categories—such as water to bridge and farmland to road, which typically involve simple backgrounds or abundant training samples—and more challenging categories, such as road to vegetation and grass to road, which are characterized by complex backgrounds, ambiguous boundaries, and limited sample sizes. Across this diverse range of semantic changes, our method achieves significant advantages in both the IoU and F1 metrics. Specifically, for the easily detectable categories—water to bridge, bridge to water, water to road, and farmland to road—our model attains IoU scores of 86.83%, 83.58%, 71.68%, and 68.40%, respectively, significantly outperforming other comparison methods. These results highlight the strong capability of our model in semantic discrimination and spatial reasoning. More importantly, our method also performs excellently on several challenging categories, including road to vegetation, road to grass, road to bare land, and grass to road, which suffer from sample scarcity and semantic ambiguity. These types of changes are often affected by similar surface textures, seasonal variability, and complex contextual information. Traditional methods are prone to false positives and missed detections under such conditions, whereas our model consistently demonstrates strong robustness and generalization ability by accurately identifying these subtle semantic transformations.
Furthermore, we evaluate the performance of the latest RS foundation model, LWGANet (Lu et al., 2025), on our proposed dataset (as shown in Table 5). The results further underscore the challenging nature of our dataset. By incorporating fine-grained and practically meaningful semantic change categories in traffic scenarios, our dataset provides a more rigorous and instructive benchmark for training and evaluating CD models in real-world applications.
Results for Binary Change Detection on RB-SCD Dataset: As shown in Table LABEL:tab:BCD, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics. In terms of the IoU metric, our model ranks first with a score of 65.83%, slightly outperforming ELGC-Net (Noman et al., 2024), which achieves 65.53%. This indicates that MFDCD demonstrates the best performance in terms of overlap between the predicted and ground truth change regions, enabling more precise localization of change areas. Regarding the F1, our MFDCD model also leads with a score of 79.39%, followed closely by ELGC-Net at 79.18%. Since the F1 is a harmonic mean of Pre and Rec, this result suggests that MFDCD achieves an excellent balance between detection accuracy and completeness. For recall, USSFC-Net (Lei et al., 2023) achieves the highest score of 75.07% among all models, while our model follows closely with a recall of 74.13%, ranking second. This demonstrates that USSFC-Net has a strong ability to detect all change areas, and our model also exhibits competitive performance in capturing changed regions. In terms of precision, our model outperforms others with a score of 85.47%, while ELGC-Net follows at 85.15%. This implies that our model has a lower false positive rate when identifying change areas. Overall, the results of the BCD task further indicate that existing methods still have room for improvement when applied to our RB-SCD dataset.
Results on Public CD Datasets: To further prove the validity of our model, we also compare our proposed MFDCD with thirteen state-of-the-art CD techniques. To ensure a fair comparison, we reproduce these methods using publicly available code with default parameters. Table 7 presents the quantitative evaluations of various methods on three public CD datasets, including IoU, F1, Rec, and Pre metrics. As observed in Table 7, our method achieves outstanding results across all three publicly available datasets. Specifically, on the WHU-CD dataset, our approach outperforms the second-best method by 2.69% and 1.5% in IoU and F1, respectively. On the SYSU-CD dataset, it achieves improvements of 4.43% and 3.05%, while on the LEVIR-CD dataset, it surpasses the second-best method by 1.06% and 0.62%. These results further demonstrate the effectiveness and robustness of our method.
Figure 7: Qualitative visualization for semantic change detection on RB-SCD dataset. (a) $T_{1}$ images. (b) $T_{2}$ images. (c) HGINet. (d) CdSC. (e) DEFO. (f) ChangeCLIP. (g) Ours. (h) Label.
Figure 8: Qualitative visualization for binary change detection on RB-SCD dataset. (a) $T_{1}$ images. (b) $T_{2}$ images. (c) BIT. (d) ChangeFormer. (e) ICIF. (f) TFI-GR. (g) DMINet. (h) USSFC-Net. (i) ELGC-Net. (j) Ours. (k) Label. The rendered colors represent true positives (white), false positives (red), true negatives (black), and false negatives (blue).
Figure 9: Heatmaps of our proposed modules. (a) and (g) Input images. (b) and (h) Low-level features from the backbone network. (c) and (i) High-level features from the backbone network. (d) and (j) Feature maps after undergoing the ASFF block. (e) and (k) Feature maps after undergoing the BTFF block. (f) and (l) Feature maps after undergoing the TFF module.
5.4 Quantitative Analysis.
As shown in Figure 7, we compared the visualized results of the proposed MFDCD with other state-of-the-art methods on RB-SCD dataset. From the figure, it can be observed that our method outperforms others in identifying small semantic change regions. For instance, as shown in the third row, our method better distinguishes between changes from grass to road and from vegetation to road—two types of changes with similar characteristics. Regarding pseudo-changes caused by background composition, such as in the fifth row, where a boat passing under a bridge with its bow color similar to the bridge’s edge color, both DEFO (Li et al., 2024) and ChangeCLIP (Dong et al., 2024) made incorrect predictions, while our method accurately detected the changes. Additionally, the proposed dataset presents significant challenges. For example, in the first row, all methods struggle to detect semantic changes under road shadows. As shown in the second row, farmland in autumn and winter appears visually similar to bare land, and in the fourth row, farmland in spring and summer shares similar colors with water, resulting in low inter-class differences. In such challenging scenes, other methods show many false detections, while our method performs slightly better but still leaves room for improvement.
When comparing the BCD results of each method with the ground truth in Figure  8, it can be observed that for the unchanged areas (black regions representing true negatives), most models show relatively accurate results, with fewer false positives. In contrast, for the changed areas, the performance varies among models. Some models exhibit noticeable missed detections (blue regions representing false negatives), indicating they fail to identify all actual changes. Many models are capable of detecting building—related changes, which are common in the dataset. For example, in areas where buildings are newly constructed or modified, most models can capture at least part of the change, though the precision of boundary detection varies. In regions where there is a transition from bare land to road, our MFDCD and some others show better performance in accurately delineating the change area. Some methods, however, have varying degrees of missed detections. For instance, methods such as BIT (Chen et al., 2022) and ChangeFormer (Bandara and Patel, 2022) seem to have less accurate delineation compared to our method and ELGC-Net (Noman et al., 2024). In areas where buildings are under construction in one image and completed in another, different models have different performances. Some models, like ICIF (Feng et al., 2022) and TFI-GR (Li et al., 2022), can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives). Other models may miss parts of the change. In areas where the change from bare land to road is affected by shadows (resulting in pseudo-changes), most models struggle. Except for our MFDCD, which shows relatively better performance in distinguishing true changes from pseudo-changes caused by shadows, other models either misclassify the shadow areas as changes (false positives in red) or miss the actual changes (false negatives in blue).
Table 8: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets. All values are reported as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Method DFC TFF RB-SCD LEVIR -CD SYSU -CD WHU -CD
(a) × × 50.45 81.65 68.98 86.65
(b) × 55.89 83.25 71.10 89.28
(c) × 55.23 82.97 70.89 88.50
(d) 57.52 84.69 72.63 90.43
Table 9: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Method RB-SCD LEVIR -CD SYSU -CD WHU -CD
(a) - 55.23 82.97 70.89 88.50
(b) $F_{i}^{1}+H_{i}^{2},i\in\left\{1,2,3,4\right\}$ 55.84 83.52 71.42 89.34
(c) $F_{i}^{t}+LL_{i}^{t},i\in\left\{1,2,3,4\right\}$ 56.10 83.67 71.68 89.45
(d) $F_{i}^{1}+H_{i}^{2},F_{i}^{2}+H_{i}^{1},i\in\left\{1,2\right\}$ 55.70 83.39 71.23 88.90
(e) $F_{i}^{t}+LL_{i}^{t},i\in\left\{3,4\right\}$ 55.86 83.51 71.30 89.24
(f) ASFF 56.42 84.02 72.06 89.98
(g) BTFF 56.87 83.90 71.85 89.77
(h) ASFF+BTFF 57.52 84.69 72.63 90.43
Table 10: Effectiveness of graph filters with different numbers $k$ in TFF. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Method $k$ RB-SCD LEVIR-CD SYSU-CD WHU-CD
(a) - 55.89 83.25 71.10 89.28
(b) 3 56.74 83.84 71.98 89.87
(c) 4 57.13 84.13 72.31 90.10
(d) 5 57.52 84.69 72.63 90.43
(e) 6 57.40 84.26 72.25 90.38
5.5 Ablation Study
To demonstrate the effectiveness of our proposed modules, we conduct extensive experiments on RB-SCD and three public datasets. As shown in Table 8, when the DFC module and TFF module are added to the baseline model, the mIoU on RB-SCD dataset increased by 5.44% and 4.78%, respectively. Additionally, the IoU on the LEVIR-CD, SYSU-CD, and WHU-CD datasets improved by 1.6%/2.12%/2.63% and 1.32%/1.91%/1.85%, respectively. As shown in Table 9, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module. Initially, we experiment with different coupling methods for ordinary features and frequency domain features. By comparing rows (b) and (c) with row (h), we can observe that ASFF and BTFF effectively dynamically couple visual features with frequency domain features, achieving the best results. Comparing rows (d) and (f), it can be seen that combining the low-level features at time $T_{1}$ with the corresponding high-frequency features at time $T_{2}$, due to semantic mismatches, results in worse performance compared to our designed ASFF. Comparing rows (e) and (g), it is clear that our designed BTFF captures the global structural changes across time, outperforming the coupling of high-level features and low-frequency features at the same time point. As shown in Table 10, we conduct ablation experiments on the number of layers in the frequency filter of TFF. As the value of $k$ increased, the model’s performance improved. However, when the number of filter layers reached six, the performance stabilized or slightly declined. Subsequent experiments are based on the optimal setting, $k=5$. To further demonstrate the effectiveness of our modules, we visualized the intermediate layer features of the network as heatmaps. As shown in Figure 9, the proposed modules effectively enhance the detection capability for roads and bridges.
6 Conclusion
In this work, we introduce RB-SCD dataset, a SCD dataset specifically designed for complex traffic scenarios. The dataset features fine-grained pixel-level annotations across 11 semantic change categories and one background class, encompassing a wide range of road and bridge types. It also covers diverse geographic regions across multiple countries and urban environments, making it become a comprehensive and challenging benchmarks to date for road and bridge change analysis in RS imagery. Notably, RB-SCD supports both BCD and SCD tasks, offering flexibility for different application scenarios. By incorporating both common and rare semantic transformations, RB-SCD provides a valuable resource for the development and evaluation of CD models in real-world, high-density traffic contexts.
To address the unique challenges presented by this dataset, we propose a novel multimodal CD framework, MFDCD, which leverages both spatial and frequency domain representations to enhance feature learning. By introducing frequency-domain analysis into the CD pipeline, MFDCD captures subtle semantic differences and structural changes that may be overlooked by conventional spatial-domain approaches. Our method also incorporates an effective fusion mechanism to integrate multimodal features, enabling more accurate detection of complex changes such as occluded road regions. Extensive experiments conducted on the RB-SCD dataset and public CD datasets demonstrate the superior performance of MFDCD across various metrics when compared to state-of-the-art CD methods. The results highlight the robustness, generalization ability, and semantic discrimination power of our approach in both straightforward and highly ambiguous change scenarios.
In future work, we plan to extend our exploration of road and bridge change detection by integrating additional modalities such as elevation data or temporal sequences, and further investigating the role of frequency-domain representations in improving the interpretability and adaptability of CD models. As the label imbalance analysis and challenges for RB-SCD dataset discussed in Section  3.2, we also plan to expand the scale of the dataset and enrich it with more detailed spatiotemporal annotations to support broader research tasks such as urban infrastructure evolution and intelligent transportation systems. Furthermore, while MFDCD achieves strong performance, it is not yet lightweight enough for deployment in resource-constrained environments. Future research will focus on model compression and efficient architecture design to improve its applicability in real-time and edge scenarios.
CRediT authorship contribution statement
Qingling Shu: Writing—review & editing, Writing—original draft, Methodology, Funding acquisition, Formal analysis, Data curation. Sibao Chen: Software, Resources, Project administration. Zhihui You: Validation, Supervision. Wei Lu: Validation, Supervision. Jin Tang: Validation, Supervision. Bin Luo: Validation, Supervision.
Declaration of competing interest
The authors declare that they have no known competing finan- cial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgement
This work was supported in part by NSFC Key Project of International (Regional) Cooperation and Exchanges (No. 61860206004), NSFC Key Project of Joint Fund for Enterprise Innovation and Development (No. U20B2068, U24A20342) and National Natural Science Foundation of China (No. 61976004).

### v2

@@
 Abstract
-Accurate detection of changes in roads and bridges, such as construction, renovation, and demolition, is essential for urban planning and traffic management.
-However, existing methods often struggle to extract fine-grained semantic change information due to the lack of high-quality annotated datasets in traffic scenarios.
-To address this, we introduce the Road and Bridge Semantic Change Detection (RB-SCD) dataset, a comprehensive benchmark comprising 260 pairs of high-resolution remote sensing images from diverse cities and countries.
-RB-SCD captures 11 types of semantic changes across varied road and bridge structures, enabling detailed structural and functional analysis.
-Building on this dataset, we propose a novel framework, Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain.
-MFDCD includes a Dynamic Frequency Coupler (DFC) that fuses hierarchical visual features with wavelet-based frequency components, and a Textual Frequency Filter (TFF) that transforms CLIP-derived textual features into the frequency domain and applies graph-based filtering.
-Experimental results on RB-SCD and three public benchmarks demonstrate the effectiveness of our approach.
-keywords:
-[…]
-1 Introduction
-In the field of remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes (Pekel et al., 2016; Song et al., 2018).
+With the rapid modernization of urban transportation, the accurate detection of changes such as construction, renovation, and demolition of roads and bridges plays a crucial role in urban planning and traffic management.
+However, existing methods often struggle to precisely extract semantic change information in such scenarios.
+A major limitation is the scarcity of fine-grained annotated change detection (CD) datasets in traffic scenes, which hinders the training and optimization of relevant algorithms.
+To address this, we introduce Road and Bridge Semantic Change Detection (RB-SCD) dataset, a comprehensive benchmark containing 260 pairs of high-resolution remote sensing images.
+The dataset spans a wide range of geographic regions and includes diverse road and bridge types across more than ten cities in multiple countries, incorporating 11 distinct categories of semantic changes, which collectively enable detailed structural and functional change analysis.
+Leveraging the rich diversity and complexity of RB-SCD, we further propose a novel framework, Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multiple characteristics of multimodal features in the frequency domain for the first time.
+Technically, MFDCD consists of two components: Dynamic Frequency Coupler (DFC) and the Textual Frequency Filter (TFF).
+The DFC deeply couples hierarchical visual features with wavelet-derived frequency components, thereby enhancing the perception of detailed and cross-temporal global structural changes.
+The TFF converts textual features extracted by the pre-trained CLIP model into the frequency domain via Fourier transform.
+Then it applies graph-based filtering to extract salient frequency components, which are fused with visual features to enable joint multimodal representation learning.
+Experimental results show that MFDCD performs well on RB-SCD and three public datasets.
+The RB-SCD dataset, with its rich semantic annotations, serves as a challenging yet valuable resource for road and bridge change analysis in traffic environments.
+The code and dataset will be available at https://github.com/DaGuangDaGuang/RB-SCD.
+Index Terms:
+Change detection (CD), road and bridge semantic change detection dataset, dynamic frequency coupler, textual frequency filter, remote sensing (RS), semantic map guidance.
+I Introduction
+In the field of remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes [1, 2].
 With the acceleration of urbanization and the continuous development of transportation infrastructure, CD in traffic scenes is crucial for traffic planning and facility management.
-In traffic planning, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades (Lin and Habib, 2022; K S et al., 2024; Wang et al., 2025).
+In traffic planning, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades [3][4][5].
 For instance, the emergence of new overpasses, flyovers, or road expansions needs to be timely integrated into transportation maps to ensure up-to-date routing and navigation services.
@@
 Roads and bridges are not only physical assets but also key elements in the digital infrastructure of intelligent transportation systems.
-However, as shown in Table  1, traditional binary change detection (BCD) datasets (Ji et al., 2018; Chen and Shi, 2020; Shi et al., 2021), etc.
+However, as shown in Tab.
+I, traditional binary change detection (BCD) datasets [6][7][8], etc.
 primarily focus on building CD and only provide simple information about whether a change has occurred, represented by 0 and 1 for no change and change, respectively.
-On the other hand, existing semantic change detection (SCD) datasets (Caye Daudt et al., 2019; Yuan et al., 2022; Yang et al., 2022a), etc.
+On the other hand, existing semantic change detection (SCD) datasets [9][10][11], etc.
 mainly focus on changes in land cover, such as building changes and vegetation changes.
 Changes related to roads and bridges in traffic scenes are either limited in number or completely absent.
-To address this issue, as annotation examples shown in Figure 1, we propose a new dataset specifically designed for SCD of roads and bridges in traffic scenes: Road and Bridge Semantic Change Detection (RB-SCD) dataset.
+Most existing road extraction datasets [12][13][14] mainly focus on binary pixel segmentation of roads and cannot be used to focus on road changes.
+To address this issue, as annotation examples shown in Fig.
+1, we propose a new dataset specifically designed for SCD of roads and bridges in traffic scenes: Road and Bridge Semantic Change Detection (RB-SCD) dataset.
 RB-SCD provides semantic annotations for 11 distinct categories of road and bridge changes, effectively addressing insufficient granularity of information and lack of specificity in detecting road and bridge changes in traffic scenes.
 In addition to the RB-SCD dataset, we have also developed a new framework to further facilitate CD performance, the Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain.
-Considering CD methods, traditional approaches often rely on convolutional neural network architectures (Caye Daudt et al., 2018; Fang et al., 2022; Feng et al., 2022), which are insufficient to fully exploit the rich spatiotemporal information in images.
-Some studies (Chen et al., 2022; Bandara and Patel, 2022; Feng et al., 2023) have incorporated transformers into the network for global modeling.
+Considering CD methods, traditional approaches often rely on convolutional neural network architectures [15][16][17], which are insufficient to fully exploit the rich spatiotemporal information in images.
+Some studies [18][19][20] have incorporated transformers into the network for global modeling.
 However, these methods still face challenges in accurately capturing and effectively distinguishing complex textures and structural changes.
 Research has also shown that relying solely on single-modal data has inherent limitations.
-Recently, some methods  (Yang et al., 2022b; Zhang et al., 2023; Dong et al., 2024) have introduced multimodal features to overcome the shortcomings of single modalities, such as combining visual and textual information or integrating lidar and hyperspectral data.
+Recently, some methods [21][22][23] have introduced multimodal features to overcome the shortcomings of single modalities, such as combining visual and textual information or integrating lidar and hyperspectral data.
 However, these approaches often fail to fully exploit the characteristics of multimodal features, focusing only on modeling the commonalities of multimodal data while neglecting the distributional differences among different modalities.
-Table 1: The public datasets for urban change detection with optical remote sensing imagery.
-Dataset Resolution (m) Image pairs Image size (pixels) Change type Objects of interest Classes
-LMR (Low-to-Medium Resolution)
-OSCD  (Daudt et al., 2018) 10 24 $600\times 600$ Binary Land cover
-S2MLTCP  (Leenstra et al., 2021) 10 1520 $600\times 600$ Binary Land cover
-ZY3  (Zhang and Shi, 2020) 5.8 1 $1154\times 740$ Binary Land cover
-Landsat-SCD  (Yuan et al., 2022) 30 8468 $416\times 416$ Semantic Land cover 4
-HR (High-Resolution)
-WV3  (Zhang and Shi, 2020) 2 2 $1431\times 1431$ Binary Land cover
-QB  (Zhang and Shi, 2020) 2.4 1 $1154\times 740$ Binary Land cover
-VHR (Very High-Resolution)
-AICD  (Bourdis et al., 2011) 0.5 1000 $800\times 600$ Binary Land cover
-LEVIR-CD  (Chen and Shi, 2020) 0.5 637 $1024\times 1024$ Binary Buildings
-WHU  (Ji et al., 2018) 0.2 1 $32207\times 15354$ Binary Buildings
-CDD  (Lebedev et al., 2018) 0.03 - 1 16000 $256\times 256$ Binary Buildings
-SYSU  (Shi et al., 2021) 0.5 16000 $256\times 256$ Binary Buildings
-BDD  (Adriano et al., 2021) 1.0 1147 $250\times 250$ Semantic Buildings 1
-HRSCD  (Caye Daudt et al., 2019) 0.5 291 $10000\times 10000$ Semantic Land cover 5
-SECOND  (Yang et al., 2022a) 0.5 - 3 4662 $512\times 512$ Semantic Land cover 6
-Hi-UCD  (Tian et al., 2020) 0.1 1293 $1024\times 1024$ Semantic Land cover 9
+TABLE I: Public remote sensing change detection and road extraction datasets. For image counting, the unit of change detection datasets is pairs, while that of road extraction is pieces/images.
+Dataset Resolution (m) Image count Image size (pixels) Type Objects of interest Classes
+Change detection
+OSCD  [24] 10 24 $600\times 600$ Binary Land cover
+S2MLTCP  [25] 10 1520 $600\times 600$ Binary Land cover
+ZY3  [26] 5.8 1 $1154\times 740$ Binary Land cover
+WV3  [26] 2 2 $1431\times 1431$ Binary Land cover
+QB  [26] 2.4 1 $1154\times 740$ Binary Land cover
+AICD  [27] 0.5 1000 $800\times 600$ Binary Land cover
+LEVIR-CD  [7] 0.5 637 $1024\times 1024$ Binary Buildings
+WHU  [6] 0.2 1 $32207\times 15354$ Binary Buildings
+CDD  [28] 0.03 - 1 16000 $256\times 256$ Binary Buildings
+SYSU  [8] 0.5 20000 $256\times 256$ Binary Buildings
+BDD  [29] 1.0 1147 $250\times 250$ Semantic Buildings 1
+Landsat-SCD  [10] 30 8468 $416\times 416$ Semantic Land cover 4
+HRSCD  [9] 0.5 291 $10000\times 10000$ Semantic Land cover 5
+SECOND  [11] 0.5 - 3 4662 $512\times 512$ Semantic Land cover 6
+Hi-UCD  [30] 0.1 1293 $1024\times 1024$ Semantic Land cover 9
+Road extraction
+CHN6-CUG  [12] 0.5 4511 $512\times 512$ Binary Road
+Massachusetts  [13] 1.5 1000 $1500\times 1500$ Binary Road
+Google Earth  [14] 1.2 224 $512\times 512$ Binary Road
 RB-SCD (Ours) 0.59 260 $270\times 534$ $7215\times 4366$ Semantic Roads and bridges 8
-Figure 1: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represents pre- and post-change images, respectively.
+Figure 1: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represent the pre- and post-change images, respectively. Label is the ground truth.
 Figure 2: Visualization of visual and textual frequency spectrum.
-In this work, as shown in Figure 2, we observe that the decomposition and utilization of features in the frequency domain remain underexplored, particularly the fine-grained change information embedded in the high-frequency visual features, the large-scale distribution information in the low-frequency features, and the frequency-domain representation of textual features.
+In this work, as shown in Fig.
+2, we observe that the decomposition and utilization of features in the frequency domain remain underexplored, particularly the fine-grained change information embedded in the high-frequency visual features, the large-scale distribution information in the low-frequency features, and the frequency-domain representation of textual features.
 The potential patterns of multimodal information in the frequency domain have not been adequately studied.
@@
 The BTFF block addresses this by combining high-level features with low-frequency features across time points, enabling the model to understand the trend of changes and global structural changes over time.
-TFF introduces a pre-trained CLIP (Radford et al., 2021) model to obtain textual information, which is further expanded to get text descriptors.
+TFF introduces a pre-trained CLIP [31] model to obtain textual information, which is further expanded to get text descriptors.
 Text features contain semantic information but lack spatial features, and discrete Fourier transform (DFT) can convert them into frequency domain information, making it easier to align with the frequency information of image features.
 By constructing graph filters to aggregate key information, the processed textual features are deeply integrated with visual features to provide multimodal support for SCD.
-Our contributions are as follows: {adjustwidth}1.5em0pt•  A comprehensive benchmark for SCD in traffic scenarios.
+Our contributions are as follows:
+  • A comprehensive benchmark for SCD in traffic scenarios.
 We construct a new dataset named RB-SCD, which focuses on SCD of roads and bridges.
@@
 With its wide geographic diversity and complex traffic infrastructures, RB-SCD supports fine-grained structural and functional change analysis for transportation environments.
-{adjustwidth}1.5em0pt•  A novel multimodal frequency-driven CD framework.
+  • A novel multimodal frequency-driven CD framework.
 We propose Multimodal Frequency-Driven Change Detector (MFDCD), the first framework to introduce frequency-domain multimodal feature integration for CD.
-It consists of two main modules: (1) Dynamic Frequency Coupler (DFC), which fuses hierarchical visual features with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception; (2) Textual Frequency Filter (TFF), which transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative multimodal representation learning.
-{adjustwidth}1.5em0pt•  Comprehensive experiments on multiple benchmarks.
+It consists of two main modules: (1) Dynamic Frequency Coupler (DFC), which fuses hierarchical visual features with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception; (2) Textual Frequency Filter (TFF), which transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning.
+  • Comprehensive experiments on multiple benchmarks.
 Extensive experiments on the proposed RB-SCD and three public CD datasets demonstrate that MFDCD achieves superior performance and generalization ability in identifying complex semantic changes in traffic scenes.
-2 Related Work
-2.1 CD Datasets and Methods
-Existing datasets, as shown in Table 1, such as  (Ji et al., 2018; Lebedev et al., 2018; Chen and Shi, 2020; Shi et al., 2021), are primarily designed for BCD tasks and cannot provide specific semantic change information.
-For SCD, datasets like  (Caye Daudt et al., 2019; Yuan et al., 2022; Yang et al., 2022a) have established some benchmarks.
+II Related Work
+II-A CD Datasets and Methods
+Existing datasets, as shown in Tab.
+I, such as  [6][7][8][28], are primarily designed for BCD tasks and cannot provide specific semantic change information.
+For SCD, datasets like  [9][10][11] have established some benchmarks.
 However, these datasets mainly focus on changes in land cover, with little attention paid to the changes in roads and bridges in traffic scenes.
@@
 Figure 3: (a) Statistical distribution of pixels for semantic change categories. (b) Train set. (c) Test set.
-Table 2: Specific subcategories in RB-SCD dataset.
+TABLE II: Specific subcategories in RB-SCD dataset.
 Categories Description
@@
 In recent years, deep learning has significantly improved the accuracy of CD through its ability to extract relevant features from raw data.
-Changemask (Zheng et al., 2022) decouples SCD into temporal semantic segmentation and BCD, and combines these two tasks into a unified encoder-transformer-decoder framework.
-MDENet (Liu et al., 2024) uses MDCM to capture object features in frequency and spatial domains across times, and MFGatt to reduce style impact, focusing on object location.
-JFRNet (Chang et al., 2024) introduces a three-branch network that takes joint bi-temporal features as unified input, enabling each branch to perceive the semantic relevance of the two time points without any additional interaction operations.
-DecoderFusion (Li et al., 2024) enhances CD decoding features by leveraging the feature differences between two semantic segmentation decoding branches.
-CdSC (Wang et al., 2024) introduces a cross-differential semantic consistency network that explores the differences in bi-temporal features while maintaining their semantic consistency.
-HGINet (Long et al., 2024) builds a multi-level perception aggregation network with a pyramid architecture to extract features that distinguish different categories across multiple levels.
+Changemask [32] decouples SCD into temporal semantic segmentation and BCD, and combines these two tasks into a unified encoder-transformer-decoder framework.
+MDENet [33] uses MDCM to capture object features in frequency and spatial domains across times, and MFGatt to reduce style impact, focusing on object location.
+JFRNet [34] introduces a three-branch network that takes joint bi-temporal features as unified input, enabling each branch to perceive the semantic relevance of the two time points without any additional interaction operations.
+DecoderFusion [35] enhances CD decoding features by leveraging the feature differences between two semantic segmentation decoding branches.
+CdSC [36] introduces a cross-differential semantic consistency network that explores the differences in bi-temporal features while maintaining their semantic consistency.
+HGINet [37] builds a multi-level perception aggregation network with a pyramid architecture to extract features that distinguish different categories across multiple levels.
 However, these methods only consider single-modal data, i.e., images, while neglecting the rich semantic information embedded in multimodal data, thus encountering a performance bottleneck.
-2.2 Multimodal Learning
+II-B Multimodal Learning
 Recent studies have shown that multimodal information is beneficial for visual image analysis.
-LAVT (Yang et al., 2022b) and VLT (Ding et al., 2023) integrate textual features into visual features through language encoders.
+LAVT [21] and VLT [38] integrate textual features into visual features through language encoders.
 Zhang et al.
-(Zhang et al., 2023) propose a convolutional neural network method with an attention mechanism to enhance feature extraction for lidar data.
+[22] propose a convolutional neural network method with an attention mechanism to enhance feature extraction for lidar data.
 Their approach leverages the elevation information from lidar data and the spatial-spectral information of hyperspectral data for multimodal feature fusion.
-MedCLIP (Koleilat et al., 2024) uses text prompts to generate segmentation for clinical scans, adopting a novel decoupled hard-negative noise contrastive estimation loss to fine-tune the BiomedCLIP model.
-ChangeCLIP (Dong et al., 2024) designs a novel CD framework that utilizes semantic information from image-text pairs and introduce an innovative differential feature compensation module to capture fine-grained semantic changes between them.
-2.3 Frequency Transform in Deep Learning
+MedCLIP [39] uses text prompts to generate segmentation for clinical scans, adopting a novel decoupled hard-negative noise contrastive estimation loss to fine-tune the BiomedCLIP model.
+ChangeCLIP [23] designs a novel CD framework that utilizes semantic information from image-text pairs and introduce an innovative differential feature compensation module to capture fine-grained semantic changes between them.
+II-C Frequency Transform in Deep Learning
 In image processing, the combination of CNN and frequency domain information has been widely applied in tasks such as image restoration, super-resolution, and semantic segmentation.
-W-Unet (Zhao et al., 2021) replaces down-sampling operations with DWT to reduce the impact of image noise while avoiding information loss.
-FFT-ReLU  (Mao et al., 2023) performs image deblurring tasks in the frequency domain based on fast Fourier transform.
-ARFFT (Zhu et al., 2023) proposes a spatial-frequency fusion block to enhance the representation capability of Transformers and extend the receptive field across the entire image.
-XNet (Zhou et al., 2023) utilizes wavelet transforms to transfer visual information to the frequency domain, leveraging high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks.
+W-Unet [40] replaces down-sampling operations with DWT to reduce the impact of image noise while avoiding information loss.
+FFT-ReLU  [41] performs image deblurring tasks in the frequency domain based on fast Fourier transform.
+ARFFT [42] proposes a spatial-frequency fusion block to enhance the representation capability of Transformers and extend the receptive field across the entire image.
+XNet [43] utilizes wavelet transforms to transfer visual information to the frequency domain, leveraging high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks.
 Given the characteristics of RS images, such as high resolution and susceptibility to spectral variations caused by seasonal changes, using only frequency domain features often results in the loss of critical spatial information.
@@
 Figure 5: Examples of road and bridge types, where the first two columns are roads, and the last three columns are bridges.
-3 RB-SCD Dataset
-3.1 Data Annotation and Analysis
-Rich Semantic Annotations: As shown in Figure 1 and Figure 3, unlike traditional BCD datasets, we have defined eight major categories (with specific subcategories detailed in Table 2 ), including road, bridge, bare land, building, farmland, water, grass and vegetation.
+III RB-SCD Dataset
+III-A Data Annotation and Analysis
+Rich Semantic Annotations: As shown in Fig.
+1 and Fig.
+3, unlike traditional BCD datasets, we have defined eight major categories (with specific subcategories detailed in Tab.
+II), including road, bridge, bare land, building, farmland, water, grass and vegetation.
 The RB-SCD dataset provides detailed semantic change annotations for 11 distinct types of semantic changes and a background category across these eight categories.
 Figure 6: Overview of our proposed Multimodal Frequency-Driven Change Detector (MFDCD).
+Firstly, the Dynamic Frequency Coupler (DFC) couples hierarchical visual features extracted by ResNet50 with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception.
+Meanwhile, the Textual Frequency Filter (TFF) transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning.
+Finally, encoder features are fed into the decoder to obtain the final semantic change detection results.
 Changes such as road widening, new road construction, and bridge construction or modification are reflected through the corresponding semantic change types, supporting in-depth exploration of the intrinsic mechanisms of traffic scenario evolution.
@@
 All annotations are carried out by a professional team in RS and transportation, and underwent multiple rounds of rigorous review and quality control to ensure annotation accuracy and semantic consistency.
-Diverse Geographic Coverage: As shown in Figure 4, following the principles in (Long et al., 2021), we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set.
+Diverse Geographic Coverage: As shown in Fig.
+4, following the principles in [44], we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set.
 The dataset integrates satellite images obtained primarily from Google Earth, covering various regions worldwide, including Chinese cities such as Hefei, Xi’an, Chongqing and Guangzhou, as well as representative areas from Japan, Singapore, Nigeria, Italy, America, and Indonesia, etc.
 This extensive geographic coverage ensures that the dataset encompasses traffic scenes under varying climatic conditions, geological structures, and urban planning contexts, thereby guaranteeing its diversity and representativeness.
-Multiple Road and Bridge Types: As shown in Figure 5, the dataset encompasses various types of roads and bridges from different geographical regions and urban environments.
+Multiple Road and Bridge Types: As shown in Fig.
+5, the dataset encompasses various types of roads and bridges from different geographical regions and urban environments.
 The road categories include concrete and asphalt roads used for urban main roads, secondary roads, residential streets, and highways, comprehensively reflecting the common transportation routes found in cities.
@@
 This richness in road and bridge types enables the dataset to represent complex traffic environments and provides robust support for training CD models across multiple scales and diverse scenarios.
-3.2 Label Imbalance Analysis and Challenges
-As shown in Figure  3, our RB-SCD dataset exhibits a certain degree of sample imbalance, primarily manifested in the unequal distribution of different road and bridge types as well as change categories.
-Due to the diversity in data collection times and geographical regions, some categories of roads, bridges, and change types have fewer samples.
-In terms of semantic changes, there is also an imbalance in the types of changes in the dataset, especially for fine-grained CD tasks.
+III-B Label Imbalance Analysis and Challenges
+As shown in Fig.
+3, our RB-SCD dataset exhibits a certain degree of sample imbalance In terms of semantic changes, there is an imbalance in the types of changes in the dataset, especially for fine-grained CD tasks.
 For some types of changes (e.g., road widening, bridge construction), there are more samples, while for other types of changes (e.g., road repair, demolition), there are fewer samples.
@@
 To improve the accuracy and robustness of the model, we will further collect more samples of different types of changes in future work and employ techniques such as data augmentation and sample weighting to alleviate the model’s difficulty in detecting rare changes.
-4 Methodology
-4.1 Overview
-As shown in Figure 6, our MFDCD mainly includes two components: (1) DFC module effectively couples high-level and low-level visual features with high-frequency and low-frequency spectral features.
+IV Methodology
+IV-A Overview
+As shown in Fig.
+6, our MFDCD mainly includes two components: (1) DFC module effectively couples high-level and low-level visual features with high-frequency and low-frequency spectral features.
 (2) TFF module enhances the text information obtained through zero-shot inference, explores the representation of text features in the frequency domain, and promotes the interaction and fusion of multimodal data.
 The details are as follows.
-4.2 DFC Module
-Given the bi-temporal images $T_{1}$ and $T_{2}$, ResNet50 (He et al., 2016) is used as the backbone network to extract the corresponding multi-scale features $F_{i}^{t}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2)\right\}$.
+IV-B DFC Module
+Given the bi-temporal images $T_{1}$ and $T_{2}$, ResNet50 [45] is used as the backbone network to extract the corresponding multi-scale features $F_{i}^{t}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2)\right\}$.
 The features from the first two layers of the network are considered as low-level features, while the features from the last two layers are considered as high-level features.
@@
 However, high-frequency features are sensitive to noise and may amplify the influence of irrelevant regions.
-As shown in the upper-right corner of Figure  6, the ASFF block applies a novel sparsity constraint to the high-frequency features, which is dynamically adjusted during training iterations.
+As shown in the upper-right corner of Fig.
+6, the ASFF block applies a novel sparsity constraint to the high-frequency features, which is dynamically adjusted during training iterations.
 By combining the sparsity-constrained high-frequency features with low-level features, both the image’s detailed information (such as edges and textures) and local structural information are preserved.
@@
 $Conv(\cdot)$ and $Up(\cdot)$ are convolution and bilinear upsampling operation, respectively.
-Bidirectional Temporal Frequency Fusion: As shown in the lower-right corner of Figure 6.
+Bidirectional Temporal Frequency Fusion: As shown in the lower-right corner of Fig.
+6.
 Take the low-frequency features of time $T_{2}$ and the high-level features of time $T_{1}$ as an example, high-level features at time $T_{1}$ typically contain the semantic hierarchy of the image, enabling the expression of global structural changes in the image, such as road widening and bridge construction.
@@
 where $i\in\left\{3,4\right\}$, $P_{i}^{t}$ is the enhanced feature that combines low-frequency information and high-level information.
-4.3 TFF Module
-As shown in Figure 6, we leverage the powerful zero-shot inference capability of the CLIP model to obtain textual information from RS images for multimodal modeling.
+IV-C TFF Module
+As shown in Fig.
+6, we leverage the powerful zero-shot inference capability of the CLIP model to obtain textual information from RS images for multimodal modeling.
 For the categories inferred by the CLIP model, we select the top five categories with the highest probability values.
-For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ," where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself.
+For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ,” where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself.
 $E_{T}$ is a pre-trained CLIP text encoder that encodes the text embedding.
@@
 We use the text-image encoder $E_{VT}$ of the CLIP model to fuse $\mathbf{E}^{f}$ and the image features.
-After subtracting the results from the bi-temporal phases, the output is fed into the decoder (Chen et al., 2024) to obtain our CD results.
-4.4 Frequency Transform Function Analysis
+After subtracting the results from the bi-temporal phases, the output is fed into the decoder [46] to obtain our CD results.
+IV-D Frequency Transform Function Analysis
 Why use discrete wavelet transform for image features and discrete Fourier transform for textual features?
 The advantage of wavelet transform lies in its ability to analyze images at different scales and locations, making it particularly suitable for multi-level, multi-scale image analysis tasks.
-As shown in Figure  2, in remote sensing image CD, wavelet transform helps to extract high-frequency features (e.g., detailed changes) and low-frequency features (e.g., overall structure), aiding in the capture of changes at different spatiotemporal scales in the image.
+As shown in Fig.
+2, in remote sensing image CD, wavelet transform helps to extract high-frequency features (e.g., detailed changes) and low-frequency features (e.g., overall structure), aiding in the capture of changes at different spatiotemporal scales in the image.
 Considering that text is typically discrete sequential data, its changes are often global or frequency-related.
 Therefore, on the other hand, the Fourier transform can effectively reveal periodic semantic patterns in the text, which contributes to identify patterns and trends in text changes.
-5 Experiments and Results
-5.1 Public CD Datasets
-In addition to using our proposed dataset, as shown in Table 3, we have employed three public CD datasets to evaluate the robustness of our model.
-The LEVIR-CD (Chen and Shi, 2020) dataset consists of 637 pairs of RS images with a resolution of 1,024 × 1,024.
+V Experiments and Results
+V-A Public CD Datasets
+In addition to using our proposed dataset, as shown in Tab.
+III, we have employed three public CD datasets to evaluate the robustness of our model.
+The LEVIR-CD [7] dataset consists of 637 pairs of RS images with a resolution of 1,024 × 1,024.
 It mainly focuses on building area changes, such as building growth.
-The WHU-CD (Ji et al., 2018) dataset comprises a pair of high-resolution RS images with a size of 32,507 × 15,354 pixels.
+The WHU-CD [6] dataset comprises a pair of high-resolution RS images with a size of 32,507 × 15,354 pixels.
 It covers areas before and after an earthquake and primarily includes building renovations and additions.
-The SYSU-CD (Shi et al., 2021) dataset is a collection of 20,000 pairs of high-resolution aerial images, each with a size of 256 × 256 pixels.
+The SYSU-CD [8] dataset is a collection of 20,000 pairs of high-resolution aerial images, each with a size of 256 × 256 pixels.
 The images are captured in the Hong Kong area and include urban construction, suburban expansion, and vegetation changes, etc.
-Table 3: Public CD datasets used in the experiments.
+TABLE III: Public CD datasets used in the experiments.
 Datasets Train/ Val/ Test Image type Space resolution
 LEVIR-CD 7120/ 1024/ 2048 Satellite images 0.5 m/pixel
-SYSU-CD 12000/ 40000/ 4000 Aerial images 0.5 m/pixel
+SYSU-CD 12000/ 4000/ 4000 Aerial images 0.5 m/pixel
 WHU-CD 5947/ 743/ 744 Aerial images 0.2 m/pixel
-Table 4: Comparison of quantitative performance for semantic change detection on RB-SCD dataset.
+TABLE IV: Comparison of quantitative performance for semantic change detection on RB-SCD dataset.
 All metrics are expressed as percentages (%).
 Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
-Class HGINet (Long et al., 2024) CdSC (Wang et al., 2024) DEFO (Li et al., 2024) ChangeCLIP (Dong et al., 2024) Ours
+Class HGINet [37] CdSC [36] DEFO [35] ChangeCLIP [23] Ours
 IoU F1 Rec Pre IoU F1 Rec Pre IoU F1 Rec Pre IoU F1 Rec Pre IoU F1 Rec Pre
@@
 mIoU 48.65 52.75 54.63 54.95 57.52
-Table 5: Performance of the RS foundation model LWGANet on RB-SCD dataset. All metrics are expressed as percentages (%).
-Class LWGANet-L1 (Lu et al., 2025) LWGANet-L2 (Lu et al., 2025)
+TABLE V: Performance of the RS foundation model LWGANet on RB-SCD dataset. All metrics are expressed as percentages (%).
+Class LWGANet-L1 [47] LWGANet-L2 [47]
 IoU F1 Pre Rec IoU F1 Pre Rec
-background 98.3 99.14 98.84 99.44 98.27 99.13 98.8 99.45
-bare land ->road 54.17 70.27 73.28 67.5 53.17 69.42 74.46 65.02
+background 98.30 99.14 98.84 99.44 98.27 99.13 98.8 99.45
+bare land ->road 54.17 70.27 73.28 67.50 53.17 69.42 74.46 65.02
 water ->bridge 87.29 93.21 92.52 93.92 85.85 92.39 90.95 93.87
@@
 road-> bare land 38.54 55.64 78.94 42.95 38.06 55.14 74.67 43.71
-road->vegetation 15.54 26.89 55.44 17.75 13.7 24.1 68.9 14.6
-vegetation->road 59.14 74.32 75.55 73.13 60.1 75.08 78.81 71.68
-building->road 42.1 59.25 81.97 46.4 42.27 59.42 76.5 48.57
+road->vegetation 15.54 26.89 55.44 17.75 13.70 24.10 68.9 14.60
+vegetation->road 59.14 74.32 75.55 73.13 60.10 75.08 78.81 71.68
+building->road 42.10 59.25 81.97 46.40 42.27 59.42 76.50 48.57
 farmland->road 69.94 82.31 86.16 78.79 71.04 83.07 83.96 82.2
@@
 road->grass 22.78 37.11 83.44 23.86 21.87 35.89 84.47 22.79
-bridge->water 75.1 85.78 89.44 82.41 80.01 88.9 89.56 88.24
+bridge->water 75.10 85.78 89.44 82.41 80.01 88.90 89.56 88.24
 mIoU 55.12 55.71
-Table 6: Comparison of quantitative performance for binary change detection on RB-SCD dataset.
+TABLE VI: Comparison of quantitative performance for binary change detection on RB-SCD dataset.
 All metrics are expressed as percentages (%).
@@
 Model IoU F1 Rec Pre
-FC-diff (Caye Daudt et al., 2018) 57.81 73.26 69.63 77.30
-FC-ef (Caye Daudt et al., 2018) 59.05 74.26 69.05 80.31
-FC-cat (Caye Daudt et al., 2018) 60.34 75.26 71.99 78.85
-SNUNet (Fang et al., 2022) 62.90 77.22 72.47 82.64
-BIT (Chen et al., 2022) 62.21 76.71 70.97 83.69
-ChangeFormer (Bandara and Patel, 2022) 61.88 76.45 69.69 84.66
-ICIF (Feng et al., 2022) 61.80 76.39 71.93 81.43
-TFI-GR (Li et al., 2022) 64.72 78.58 73.78 84.05
-DMINet (Feng et al., 2023) 62.99 77.29 72.14 83.24
-USSFC-Net (Lei et al., 2023) 62.55 76.96 75.07 78.95
-ELGC-Net (Noman et al., 2024) 65.53 79.18 73.99 85.15
+FC-diff [15] 57.81 73.26 69.63 77.30
+FC-ef [15] 59.05 74.26 69.05 80.31
+FC-cat [15] 60.34 75.26 71.99 78.85
+SNUNet [16] 62.90 77.22 72.47 82.64
+BIT [18] 62.21 76.71 70.97 83.69
+ChangeFormer [19] 61.88 76.45 69.69 84.66
+ICIF [17] 61.80 76.39 71.93 81.43
+TFI-GR [48] 64.72 78.58 73.78 84.05
+DMINet [20] 62.99 77.29 72.14 83.24
+USSFC-Net [49] 62.55 76.96 75.07 78.95
+ELGC-Net [50] 65.53 79.18 73.99 85.15
 Ours 65.83 79.39 74.13 85.47
-Table 7: Comparison of quantitative performance of various CD techniques on public datasets.
+TABLE VII: Comparison of quantitative performance of various CD techniques on public datasets.
 All metrics are expressed as percentages (%).
@@
 Results marked with * are obtained from the original papers.
-Model LEVIR-CD (Chen and Shi, 2020) SYSU-CD (Shi et al., 2021) WHU-CD (Ji et al., 2018)
+Model LEVIR-CD [7] SYSU-CD [8] WHU-CD [6]
 IoU  F1  Rec  Pre IoU  F1  Rec  Pre IoU  F1  Rec  Pre
-FC-diff (Caye Daudt et al., 2018) 75.18 85.83 84.59 87.12 42.05 59.20 46.72 80.80 69.48 81.99 78.90 89.14
-FC-ef (Caye Daudt et al., 2018) 71.80 83.59 82.27 84.94 55.83 71.65 65.41 79.22 58.53 71.65 65.41 79.22
-FC-cat (Caye Daudt et al., 2018) 76.98 86.99 85.64 88.38 60.68 75.53 71.38 80.21 72.24 83.87 84.22 83.53
-SNUNet (Fang et al., 2022) 82.20 90.23 88.97 91.54 66.73 80.04 79.79 80.30 76.02 86.38 88.95 83.95
-BIT (Chen et al., 2022) 82.06 90.15 88.98 91.34 62.50 76.93 72.50 81.93 80.40 89.13 89.69 88.58
-ChangeFormer (Bandara and Patel, 2022) 81.93 90.07 87.38 92.92 60.46 75.36 71.06 80.22 83.88 91.23 90.76 91.71
-ICIF (Feng et al., 2022) 79.86 88.80 86.85 90.84 61.38 76.07 73.31 91.94 78.27 87.81 84.06 91.96
-TFI-GR (Li et al., 2022) 83.03 90.73 89.53 91.95 71.38 83.30 81.16 85.56 88.55 93.92 92.65 95.24
-DMINet (Feng et al., 2023) 83.14 90.79 88.89 92.78 68.20 81.10 76.82 85.88 79.53 88.60 91.24 86.10
-USSFC-Net (Lei et al., 2023) 82.54 90.44 91.82 89.09 65.75 79.34 81.52 77.27 83.92 91.26 93.20 89.40
-M-Swin * (Pan et al., 2024) 83.57 91.05 90.05 92.07 -   -   -   - 86.09 92.94 91.37 94.56
-ELGC-Net (Noman et al., 2024) 83.52 91.02 89.78 92.30 65.34 79.03 77.26 80.89 75.56 86.08 82.15 90.40
-MDENet * (Liu et al., 2024) 83.63 91.0992.68 89.55 -   -   -   - 87.74 93.47 95.49 91.54
+FC-diff [15] 75.18 85.83 84.59 87.12 42.05 59.20 46.72 80.80 69.48 81.99 78.90 89.14
+FC-ef [15] 71.80 83.59 82.27 84.94 55.83 71.65 65.41 79.22 58.53 71.65 65.41 79.22
+FC-cat [15] 76.98 86.99 85.64 88.38 60.68 75.53 71.38 80.21 72.24 83.87 84.22 83.53
+SNUNet [16] 82.20 90.23 88.97 91.54 66.73 80.04 79.79 80.30 76.02 86.38 88.95 83.95
+BIT [18] 82.06 90.15 88.98 91.34 62.50 76.93 72.50 81.93 80.40 89.13 89.69 88.58
+ChangeFormer [19] 81.93 90.07 87.38 92.92 60.46 75.36 71.06 80.22 83.88 91.23 90.76 91.71
+ICIF [17] 79.86 88.80 86.85 90.84 61.38 76.07 73.31 91.94 78.27 87.81 84.06 91.96
+TFI-GR [48] 83.03 90.73 89.53 91.95 71.38 83.30 81.16 85.56 88.55 93.92 92.65 95.24
+DMINet [20] 83.14 90.79 88.89 92.78 68.20 81.10 76.82 85.88 79.53 88.60 91.24 86.10
+USSFC-Net [49] 82.54 90.44 91.82 89.09 65.75 79.34 81.52 77.27 83.92 91.26 93.20 89.40
+M-Swin * [51] 83.57 91.05 90.05 92.07 -   -   -   - 86.09 92.94 91.37 94.56
+ELGC-Net [50] 83.52 91.02 89.78 92.30 65.34 79.03 77.26 80.89 75.56 86.08 82.15 90.40
+MDENet * [33] 83.63 91.0992.68 89.55 -   -   -   - 87.74 93.47 95.49 91.54
 Ours 84.69 91.7192.31 91.11 72.63 84.15 85.60 82.74 90.43 94.97 96.44 93.55
-5.2 Implementation Details
+V-B Implementation Details
 Our experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU.
-The AdamW (Kingma and Ba, 2014) is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01.
+The AdamW [52] is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01.
 Cross entropy loss is used as our loss function, and the maximum number of iterations is set to 40,000.
@@
 where TP, TN, FP, and FN denote the counts of true positives, true negatives, false positives, and false negatives, respectively.
-5.3 Quantitative Assessment Results
-Results for Semantic Change Detection on RB-SCD Dataset: Table  4 presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics.
+V-C Quantitative Assessment Results
+Results for Semantic Change Detection on RB-SCD Dataset: Tab.
+IV presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics.
 As shown in the table, we have defined 11 types of semantic changes along with one background category, aiming to comprehensively cover the diverse real-world scenarios of road and bridge changes.
@@
 Traditional methods are prone to false positives and missed detections under such conditions, whereas our model consistently demonstrates strong robustness and generalization ability by accurately identifying these subtle semantic transformations.
-Furthermore, we evaluate the performance of the latest RS foundation model, LWGANet (Lu et al., 2025), on our proposed dataset (as shown in Table 5).
+Furthermore, we evaluate the performance of the latest RS foundation model, LWGANet [47], on our proposed dataset (as shown in Tab.
+V).
 The results further underscore the challenging nature of our dataset.
 By incorporating fine-grained and practically meaningful semantic change categories in traffic scenarios, our dataset provides a more rigorous and instructive benchmark for training and evaluating CD models in real-world applications.
-Results for Binary Change Detection on RB-SCD Dataset: As shown in Table LABEL:tab:BCD, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics.
-In terms of the IoU metric, our model ranks first with a score of 65.83%, slightly outperforming ELGC-Net (Noman et al., 2024), which achieves 65.53%.
+Results for Binary Change Detection on RB-SCD Dataset: As shown in Tab.
+VI, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics.
+In terms of the IoU metric, our model ranks first with a score of 65.83%, slightly outperforming ELGC-Net [50], which achieves 65.53%.
 This indicates that MFDCD demonstrates the best performance in terms of overlap between the predicted and ground truth change regions, enabling more precise localization of change areas.
@@
 Since the F1 is a harmonic mean of Pre and Rec, this result suggests that MFDCD achieves an excellent balance between detection accuracy and completeness.
-For recall, USSFC-Net (Lei et al., 2023) achieves the highest score of 75.07% among all models, while our model follows closely with a recall of 74.13%, ranking second.
+For recall, USSFC-Net [49] achieves the highest score of 75.07% among all models, while our model follows closely with a recall of 74.13%, ranking second.
 This demonstrates that USSFC-Net has a strong ability to detect all change areas, and our model also exhibits competitive performance in capturing changed regions.
@@
 To ensure a fair comparison, we reproduce these methods using publicly available code with default parameters.
-Table 7 presents the quantitative evaluations of various methods on three public CD datasets, including IoU, F1, Rec, and Pre metrics.
-As observed in Table 7, our method achieves outstanding results across all three publicly available datasets.
+Tab.
+VII presents the quantitative evaluations of various methods on three public CD datasets, including IoU, F1, Rec, and Pre metrics.
+As observed in Tab.
+VII, our method achieves outstanding results across all three publicly available datasets.
 Specifically, on the WHU-CD dataset, our approach outperforms the second-best method by 2.69% and 1.5% in IoU and F1, respectively.
@@
 (f) and (l) Feature maps after undergoing the TFF module.
-5.4 Quantitative Analysis.
-As shown in Figure 7, we compared the visualized results of the proposed MFDCD with other state-of-the-art methods on RB-SCD dataset.
+V-D Quantitative Analysis.
+As shown in Fig.
+7, we compared the visualized results of the proposed MFDCD with other state-of-the-art methods on RB-SCD dataset.
 From the figure, it can be observed that our method outperforms others in identifying small semantic change regions.
 For instance, as shown in the third row, our method better distinguishes between changes from grass to road and from vegetation to road—two types of changes with similar characteristics.
-Regarding pseudo-changes caused by background composition, such as in the fifth row, where a boat passing under a bridge with its bow color similar to the bridge’s edge color, both DEFO (Li et al., 2024) and ChangeCLIP (Dong et al., 2024) made incorrect predictions, while our method accurately detected the changes.
+Regarding pseudo-changes caused by background composition, such as in the fifth row, where a boat passing under a bridge with its bow color similar to the bridge’s edge color, both DEFO [35] and ChangeCLIP [23] made incorrect predictions, while our method accurately detected the changes.
 Additionally, the proposed dataset presents significant challenges.
@@
 In such challenging scenes, other methods show many false detections, while our method performs slightly better but still leaves room for improvement.
-When comparing the BCD results of each method with the ground truth in Figure  8, it can be observed that for the unchanged areas (black regions representing true negatives), most models show relatively accurate results, with fewer false positives.
+When comparing the BCD results of each method with the ground truth in Fig.
+8, it can be observed that for the unchanged areas (black regions representing true negatives), most models show relatively accurate results, with fewer false positives.
 In contrast, for the changed areas, the performance varies among models.
@@
 Some methods, however, have varying degrees of missed detections.
-For instance, methods such as BIT (Chen et al., 2022) and ChangeFormer (Bandara and Patel, 2022) seem to have less accurate delineation compared to our method and ELGC-Net (Noman et al., 2024).
+For instance, methods such as BIT [18] and ChangeFormer [19] seem to have less accurate delineation compared to our method and ELGC-Net [50].
 In areas where buildings are under construction in one image and completed in another, different models have different performances.
-Some models, like ICIF (Feng et al., 2022) and TFI-GR (Li et al., 2022), can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives).
+Some models, like ICIF [17] and TFI-GR [48], can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives).
 Other models may miss parts of the change.
@@
 Except for our MFDCD, which shows relatively better performance in distinguishing true changes from pseudo-changes caused by shadows, other models either misclassify the shadow areas as changes (false positives in red) or miss the actual changes (false negatives in blue).
-Table 8: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets.
+TABLE VIII: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets.
 All values are reported as percentages (%).
@@
 (d) 57.52 84.69 72.63 90.43
-Table 9: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets.
+TABLE IX: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets.
 All values are reported in percentage (%).
@@
 (h) ASFF+BTFF 57.52 84.69 72.63 90.43
-Table 10: Effectiveness of graph filters with different numbers $k$ in TFF.
+TABLE X: Effectiveness of graph filters with different numbers $k$ in TFF.
 All values are reported in percentage (%).
@@
 (e) 6 57.40 84.26 72.25 90.38
-5.5 Ablation Study
+TABLE XI: Effectiveness of MFDCD with Different Backbones when Trained on Four CD Datasets.
+All Values are Reported in Percentage (%).
+Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
+Backbone Params (M) FLOPs (G) RB -SCD LEVIR -CD SYSU -CD WHU -CD
+ResNet18 13.75 11.32 55.53 83.18 71.35 89.04
+ResNet34 22.80 16.54 56.31 83.78 72.10 89.73
+ResNet50 103.78 49.02 57.52 84.69 72.63 90.43
+V-E Ablation Study
 To demonstrate the effectiveness of our proposed modules, we conduct extensive experiments on RB-SCD and three public datasets.
-As shown in Table 8, when the DFC module and TFF module are added to the baseline model, the mIoU on RB-SCD dataset increased by 5.44% and 4.78%, respectively.
+As shown in Tab.
+VIII, when the DFC module and TFF module are added to the baseline model, the mIoU on RB-SCD dataset increased by 5.44% and 4.78%, respectively.
 Additionally, the IoU on the LEVIR-CD, SYSU-CD, and WHU-CD datasets improved by 1.6%/2.12%/2.63% and 1.32%/1.91%/1.85%, respectively.
-As shown in Table 9, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module.
+As shown in Tab.
+IX, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module.
 Initially, we experiment with different coupling methods for ordinary features and frequency domain features.
@@
 Comparing rows (e) and (g), it is clear that our designed BTFF captures the global structural changes across time, outperforming the coupling of high-level features and low-frequency features at the same time point.
-As shown in Table 10, we conduct ablation experiments on the number of layers in the frequency filter of TFF.
+As shown in Tab.
+X, we conduct ablation experiments on the number of layers in the frequency filter of TFF.
 As the value of $k$ increased, the model’s performance improved.
@@
 Subsequent experiments are based on the optimal setting, $k=5$.
-To further demonstrate the effectiveness of our modules, we visualized the intermediate layer features of the network as heatmaps.
-As shown in Figure 9, the proposed modules effectively enhance the detection capability for roads and bridges.
-6 Conclusion
+To investigate the impact of backbone networks on the performance of the MFDD model in the CD tasks, as shown in Tab.
+XI, in terms of parameters and computational complexity, ResNet18 has the lowest cost (13.75M parameters, 11.32G FLOPs), while ResNet50 has the highest (103.78M parameters, 49.02G FLOPs).
+In terms of detection accuracy, for all four datasets, the deeper the backbone network, the higher the accuracy.
+However, a deeper network also leads to a doubling of parameters and computational complexity.
+In this paper, ResNet50 is selected as the backbone network.
+To more clearly illustrate the influence of our proposed modules, we visualized heatmaps for these modules.
+As depicted in columns (b)/(h) and (c)/(i) of Fig.
+9, the low-level and high-level features from the backbone network typically exhibit certain characteristics.
+The low-level features may contain fine-grained but scattered information, while the high-level features, though more abstract, might lose some detailed spatial information.
+In columns (d)/(j), (e)/(k), and (f)/(l), after undergoing the ASFF block, BTTF block and TFF module respectively, the feature maps show progressive changes.
+The ASFF block helps to fuse features in a reasonable way, making the information more integrated.
+The BTTF block further refines the features, enhancing the relevant change information.
+The TFF module finally outputs feature maps that can better highlight the change regions, eliminating some irrelevant noise and enabling the acquisition of more distinct change-related information, which showcases the efficacy of the introduced modules in feature processing for CD tasks.
+VI Conclusion
 In this work, we introduce RB-SCD dataset, a SCD dataset specifically designed for complex traffic scenarios.
@@
 Notably, RB-SCD supports both BCD and SCD tasks, offering flexibility for different application scenarios.
-By incorporating both common and rare semantic transformations, RB-SCD provides a valuable resource for the development and evaluation of CD models in real-world, high-density traffic contexts.
+By incorporating both common and rare semantic transformations, RB-SCD provides a valuable resource for the development and evaluation of CD models in traffic contexts.
 To address the unique challenges presented by this dataset, we propose a novel multimodal CD framework, MFDCD, which leverages both spatial and frequency domain representations to enhance feature learning.
@@
 In future work, we plan to extend our exploration of road and bridge change detection by integrating additional modalities such as elevation data or temporal sequences, and further investigating the role of frequency-domain representations in improving the interpretability and adaptability of CD models.
-As the label imbalance analysis and challenges for RB-SCD dataset discussed in Section  3.2, we also plan to expand the scale of the dataset and enrich it with more detailed spatiotemporal annotations to support broader research tasks such as urban infrastructure evolution and intelligent transportation systems.
+As the label imbalance analysis and challenges for RB-SCD dataset discussed in Sec.
+III-B, we also plan to expand the scale of the dataset and enrich it with more detailed spatiotemporal annotations to support broader research tasks such as urban infrastructure evolution and intelligent transportation systems.
 Furthermore, while MFDCD achieves strong performance, it is not yet lightweight enough for deployment in resource-constrained environments.
 Future research will focus on model compression and efficient architecture design to improve its applicability in real-time and edge scenarios.
-CRediT authorship contribution statement
-Qingling Shu: Writing—review & editing, Writing—original draft, Methodology, Funding acquisition, Formal analysis, Data curation.
-Sibao Chen: Software, Resources, Project administration.
-Zhihui You: Validation, Supervision.
-Wei Lu: Validation, Supervision.
-Jin Tang: Validation, Supervision.
-Bin Luo: Validation, Supervision.
-Declaration of competing interest
-The authors declare that they have no known competing finan- cial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Acknowledgement
-This work was supported in part by NSFC Key Project of International (Regional) Cooperation and Exchanges (No.
-61860206004), NSFC Key Project of Joint Fund for Enterprise Innovation and Development (No.
-U20B2068, U24A20342) and National Natural Science Foundation of China (No.
-61976004).

### v3

@@
-RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes
+Semantic Change Detection of Roads and Bridges: A Fine-grained Dataset and Multimodal Frequency-driven Detector
 Abstract
-With the rapid modernization of urban transportation, the accurate detection of changes such as construction, renovation, and demolition of roads and bridges plays a crucial role in urban planning and traffic management.
-However, existing methods often struggle to precisely extract semantic change information in such scenarios.
-A major limitation is the scarcity of fine-grained annotated change detection (CD) datasets in traffic scenes, which hinders the training and optimization of relevant algorithms.
-To address this, we introduce Road and Bridge Semantic Change Detection (RB-SCD) dataset, a comprehensive benchmark containing 260 pairs of high-resolution remote sensing images.
-The dataset spans a wide range of geographic regions and includes diverse road and bridge types across more than ten cities in multiple countries, incorporating 11 distinct categories of semantic changes, which collectively enable detailed structural and functional change analysis.
-Leveraging the rich diversity and complexity of RB-SCD, we further propose a novel framework, Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multiple characteristics of multimodal features in the frequency domain for the first time.
-Technically, MFDCD consists of two components: Dynamic Frequency Coupler (DFC) and the Textual Frequency Filter (TFF).
-The DFC deeply couples hierarchical visual features with wavelet-derived frequency components, thereby enhancing the perception of detailed and cross-temporal global structural changes.
-The TFF converts textual features extracted by the pre-trained CLIP model into the frequency domain via Fourier transform.
-Then it applies graph-based filtering to extract salient frequency components, which are fused with visual features to enable joint multimodal representation learning.
-Experimental results show that MFDCD performs well on RB-SCD and three public datasets.
-The RB-SCD dataset, with its rich semantic annotations, serves as a challenging yet valuable resource for road and bridge change analysis in traffic environments.
-The code and dataset will be available at https://github.com/DaGuangDaGuang/RB-SCD.
-Index Terms:
-Change detection (CD), road and bridge semantic change detection dataset, dynamic frequency coupler, textual frequency filter, remote sensing (RS), semantic map guidance.
-I Introduction
-In the field of remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes [1, 2].
-With the acceleration of urbanization and the continuous development of transportation infrastructure, CD in traffic scenes is crucial for traffic planning and facility management.
-In traffic planning, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades [3][4][5].
-For instance, the emergence of new overpasses, flyovers, or road expansions needs to be timely integrated into transportation maps to ensure up-to-date routing and navigation services.
-Moreover, detecting illegal road constructions or temporary blockages is also critical for enforcing urban regulations and minimizing congestion.
-From the perspective of traffic facility management, road and bridge CD enables timely identification of structural deterioration, unauthorized demolitions, or unreported construction activities.
-This semantic-level understanding is crucial for preventive maintenance, disaster response (e.g., after floods or earthquakes), and resource allocation.
-Additionally, as smart cities and autonomous transportation systems increasingly rely on accurate and up-to-date digital maps, the need for automated and fine-grained CD becomes even more pressing.
-Roads and bridges are not only physical assets but also key elements in the digital infrastructure of intelligent transportation systems.
+Accurate detection of road and bridge changes is crucial for urban planning and transportation management, yet presents unique challenges for general change detection (CD).
+Key difficulties arise from maintaining the continuity of roads and bridges as linear structures and disambiguating visually similar land covers (e.g., road construction vs.
+bare land).
+Existing spatial-domain models struggle with these issues, further hindered by the lack of specialized, semantically rich datasets.
+To fill these gaps, we introduce the Road and Bridge Semantic Change Detection (RB-SCD) dataset.
+As the first benchmark to systematically target semantic change detection of roads and bridges, RB-SCD offers comprehensive fine-grained annotations for 11 semantic change categories.
+This enables a detailed analysis of traffic infrastructure evolution.
+Building on this, we propose a novel framework, the Multimodal Frequency-Driven Change Detector (MFDCD).
+MFDCD integrates multimodal features in the frequency domain through two key components: (1) the Dynamic Frequency Coupler (DFC), which leverages wavelet transform to decompose visual features, enabling it to robustly model the continuity of linear transitions; and (2) the Textual Frequency Filter (TFF), which encodes semantic priors into frequency-domain graphs and applies filter banks to align them with visual features, resolving semantic ambiguities.
+Experiments demonstrate the state-of-the-art performance of MFDCD on RB-SCD and three public CD datasets.
+The code will be available at https://github.com/DaGuangDaGuang/RB-SCD.
+keywords:
+Road and bridge change detection, Dynamic frequency coupler, Textual frequency filter, Remote sensing
+1 INTRODUCTION
+In the field remote sensing (RS), change detection (CD) has become an essential task for monitoring environmental, urban, and infrastructure changes [1] .
+With the acceleration of urbanization and the continuous development of transportation infrastructure, accurately identifying structural and functional changes in roads and bridges helps urban authorities to assess the current state of transportation networks, forecast future demands, and make data-driven decisions about expansions, detours, and upgrades.
+Moreover, road and bridge CD enables timely identification of structural deterioration, unauthorized demolitions, or unreported construction activities [2].
 However, as shown in Tab.
-I, traditional binary change detection (BCD) datasets [6][7][8], etc.
-primarily focus on building CD and only provide simple information about whether a change has occurred, represented by 0 and 1 for no change and change, respectively.
-On the other hand, existing semantic change detection (SCD) datasets [9][10][11], etc.
-mainly focus on changes in land cover, such as building changes and vegetation changes.
+1, traditional binary change detection (BCD) datasets such as LEVIR [3], WHU [4], CDD [5], SYSU [6], merely indicate whether a change has occurred through a simple binary mask (0 for no change, 1 for change).
+They lack fine-grained semantic details.
+While existing semantic change detection (SCD) datasets like Landsat [7], HRSCD [8], SECOND [9] mainly focus on changes in land cover.
 Changes related to roads and bridges in traffic scenes are either limited in number or completely absent.
-Most existing road extraction datasets [12][13][14] mainly focus on binary pixel segmentation of roads and cannot be used to focus on road changes.
-To address this issue, as annotation examples shown in Fig.
-1, we propose a new dataset specifically designed for SCD of roads and bridges in traffic scenes: Road and Bridge Semantic Change Detection (RB-SCD) dataset.
-RB-SCD provides semantic annotations for 11 distinct categories of road and bridge changes, effectively addressing insufficient granularity of information and lack of specificity in detecting road and bridge changes in traffic scenes.
-In addition to the RB-SCD dataset, we have also developed a new framework to further facilitate CD performance, the Multimodal Frequency-Driven Change Detector (MFDCD), which integrates multimodal features in the frequency domain.
-Considering CD methods, traditional approaches often rely on convolutional neural network architectures [15][16][17], which are insufficient to fully exploit the rich spatiotemporal information in images.
-Some studies [18][19][20] have incorporated transformers into the network for global modeling.
-However, these methods still face challenges in accurately capturing and effectively distinguishing complex textures and structural changes.
-Research has also shown that relying solely on single-modal data has inherent limitations.
-Recently, some methods [21][22][23] have introduced multimodal features to overcome the shortcomings of single modalities, such as combining visual and textual information or integrating lidar and hyperspectral data.
-However, these approaches often fail to fully exploit the characteristics of multimodal features, focusing only on modeling the commonalities of multimodal data while neglecting the distributional differences among different modalities.
-TABLE I: Public remote sensing change detection and road extraction datasets. For image counting, the unit of change detection datasets is pairs, while that of road extraction is pieces/images.
+Although existing road extraction datasets [10][11] mainly focus on binary pixel segmentation of roads, they cannot be used to focus on road changes.
+Methodologically, detecting semantic changes in roads and bridges presents unique challenges that are not fully addressed by general-purpose CD methods.
+Two key difficulties stand out.
+First, unlike the discrete, monolithic changes of buildings common in datasets (first row of Fig.
+1), road networks evolve as winding linear structures and their changes often involve complex transitions from multiple, heterogeneous land-cover types such as water, vegetation, and bare land simultaneously (second row of Fig.
+1).
+Table 1: Comparison of public remote sensing change detection and road extraction datasets.
+For image counting, the unit for change detection datasets is pairs of images, whereas for road extraction datasets, it is individual image.
+Notably, no existing dataset systematically addresses road and bridge semantic changes.
 Dataset Resolution (m) Image count Image size (pixels) Type Objects of interest Classes
 Change detection
-OSCD  [24] 10 24 $600\times 600$ Binary Land cover
-S2MLTCP  [25] 10 1520 $600\times 600$ Binary Land cover
-ZY3  [26] 5.8 1 $1154\times 740$ Binary Land cover
-WV3  [26] 2 2 $1431\times 1431$ Binary Land cover
-QB  [26] 2.4 1 $1154\times 740$ Binary Land cover
-AICD  [27] 0.5 1000 $800\times 600$ Binary Land cover
 LEVIR-CD  [3] 0.5 637 $1024\times 1024$ Binary Buildings
@@
 SYSU  [6] 0.5 20000 $256\times 256$ Binary Buildings
-BDD  [29] 1.0 1147 $250\times 250$ Semantic Buildings 1
 Landsat-SCD  [7] 30 8468 $416\times 416$ Semantic Land cover 4
@@
 Massachusetts  [11] 1.5 1000 $1500\times 1500$ Binary Road
-Google Earth  [14] 1.2 224 $512\times 512$ Binary Road
-RB-SCD (Ours) 0.59 260 $270\times 534$ $7215\times 4366$ Semantic Roads and bridges 8
-Figure 1: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represent the pre- and post-change images, respectively. Label is the ground truth.
-Figure 2: Visualization of visual and textual frequency spectrum.
-In this work, as shown in Fig.
-2, we observe that the decomposition and utilization of features in the frequency domain remain underexplored, particularly the fine-grained change information embedded in the high-frequency visual features, the large-scale distribution information in the low-frequency features, and the frequency-domain representation of textual features.
-The potential patterns of multimodal information in the frequency domain have not been adequately studied.
-Therefore, we make the first attempt to develop a new architecture that maps image and textual features into different types of frequency domains based on their characteristics, enabling multimodal frequency-domain modeling for CD.
-Our MFDCD consists of two main components: Dynamic Frequency Coupler (DFC) and the Textual Frequency Filter (TFF).
-DFC decomposes bi-temporal features into high- and low-frequency components using discrete wavelet transform (DWT), and introduces an Adaptive Sparse Frequency Fusion (ASFF) block for the combination of high-frequency information and low-level features.
-During the early stages of training, ASFF enhances the sparsity of high-frequency features, focusing on significant change areas.
-In the later stages, the sparsity constraint is reduced to enhance the expressive power of the model, ultimately increasing the attention to key change areas while suppressing high-frequency noise interference.
-In this way, edge and texture information in low-level features can not only be gradually captured through convolution operations but also dynamically weighted based on the localized changes in refined high-frequency information.
-For low-frequency features, Bidirectional Temporal Frequency Fusion (BTFF) block is introduced.
-Since CD involves not only capturing local changes but also understanding the relative changes between different time points, simply combining the high-level features at a single time point with low-frequency features could cause the model to overly rely on the local structural information at that moment, neglecting the cross-temporal changes.
-The BTFF block addresses this by combining high-level features with low-frequency features across time points, enabling the model to understand the trend of changes and global structural changes over time.
-TFF introduces a pre-trained CLIP [31] model to obtain textual information, which is further expanded to get text descriptors.
-Text features contain semantic information but lack spatial features, and discrete Fourier transform (DFT) can convert them into frequency domain information, making it easier to align with the frequency information of image features.
-By constructing graph filters to aggregate key information, the processed textual features are deeply integrated with visual features to provide multimodal support for SCD.
+RB-SCD (Ours) 0.59 260 $270\times 534$ - $7215\times 4366$ Semantic Roads and bridges 8
+Figure 1: Visual comparison of change characteristics across prominent CD datasets.
+(a) LEVIR-CD and (b) SYSU-CD primarily feature discrete, monolithic building changes, (c) SECOND often focus on land-cover transitions.
+(d) RB-SCD introduces the unique challenges of detecting linear and heterogeneous change (source of change is heterogeneous) transitions, where winding roads are formed from multiple land-cover types.
+Furthermore, it presents high semantic ambiguity, with visually similar yet semantically distinct transitions such as water->road and vegetation->road, which are difficult for conventional models to distinguish.
+The limited receptive fields of recent methods like SUNNet [13], ICIF [14], USSFCNet [15] prevent them from perceiving the overall structure of road networks.
+Consequently, while they may correctly identify numerous localized patches of change, they ultimately produce fragmented and disconnected change maps, failing to assemble these patches into a single, continuous linear feature.
+Standard attention models such as BIT [16], Changeformer [17], DMINet [18], ELGCNet [19] leverage global attention to overcome this locality, their standard mechanisms treat all spatial relationships equally, lacking the explicit inductive bias required to model heterogeneous, linear structures.
+Second, the task is plagued by high semantic ambiguity.
+Visual features alone are often insufficient to resolve ambiguities, such as distinguishing a road under construction from bare land (the first and third sample of Fig.
+1 (d)), or differentiating shallow water from adjacent vegetation (second sample of Fig.
+1 (d)).
+Even multimodal models like ChangeCLIP [20], which leverage vision-language for CD, may struggle when visual features themselves are highly ambiguous, as their fusion process operates within the noisy spatial domain.
+Figure 2: Frequency-domain transformation of visual and textual data.
+(a) The DWT separates a visual scene into high-frequency (details) and low-frequency (structure) components.
+(b) The DFT converts a textual description into a global frequency spectrum, serving as a robust semantic signature for fusion.
+The limitations of existing spatial-domain methods highlight the need for a new approach that operates in a more suitable domain.
+To robustly model linear structures, the core insight is that their essential continuity of structure is naturally encoded in low-frequency components, while sharp transition boundaries correspond to high-frequency components.
+This intrinsic property makes the frequency domain a logical choice.
+As shown in Fig.
+2(a), we leverage the discrete wavelet transform (DWT) to decompose visual features and explicitly isolate these essential structural forms from fine-grained boundary details.
+To resolve deep semantic ambiguity, multimodal integration is required.
+However, prior methods suffer from fusing textual descriptions directly with unstable pixel-level features in the spatial domain.
+We instead transform text into its frequency spectrum via discrete Fourier transform (DFT) (Fig.
+2(b)).
+This frequency-domain representation naturally filters out minor textual variations, providing a stable basis for a cleaner alignment with visual features and overcoming the core limitations of spatial-level fusion.
+Based on these insights, we first introduce the Road and Bridge Semantic Change Detection (RB-SCD) dataset, a new benchmark specifically designed for SCD of roads and bridges.
+RB-SCD provides fine-grained semantic annotations for 11 distinct change categories across diverse geographic regions, enabling a detail evaluation of model capabilities in traffic domain.
+Second, we propose a novel Multimodal Frequency-Driven Change Detector (MFDCD).
+Our framework tackles the core challenges by operating in the frequency domain for both visual and textual features.
+It consists of two specialized components: (1) the Dynamic Frequency Coupler (DFC), which utilizes DWT operation to separate low-frequency components (encoding structural continuity) from high-frequency details (capturing transition boundaries).
+Within the DFC, an Adaptive Sparse Frequency Fusion (ASFF) block delineates boundaries by dynamically combining high-frequency information with low-level visual features.
+A Bidirectional Temporal Frequency Fusion (BTFF) block fuses low-level frequency and high-level visual features across time points to model the linear structure of change regions, preventing fragmentation; and (2) the Textual Frequency Filter (TFF), which first converts textual features into the frequency domain, a graph-based filtering mechanism then acts as a semantic reasoning to extract salient semantic components and joint multimodal representation learning with visual features.
+The resulting unambiguous features then guide the model toward confident semantic predictions.
 Our contributions are as follows:
-  • A comprehensive benchmark for SCD in traffic scenarios.
-We construct a new dataset named RB-SCD, which focuses on SCD of roads and bridges.
-It contains 260 large-scale high-resolution RS image pairs, covering over ten cities across multiple countries and annotating 11 typical semantic change types.
-With its wide geographic diversity and complex traffic infrastructures, RB-SCD supports fine-grained structural and functional change analysis for transportation environments.
-  • A novel multimodal frequency-driven CD framework.
-We propose Multimodal Frequency-Driven Change Detector (MFDCD), the first framework to introduce frequency-domain multimodal feature integration for CD.
-It consists of two main modules: (1) Dynamic Frequency Coupler (DFC), which fuses hierarchical visual features with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception; (2) Textual Frequency Filter (TFF), which transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning.
-  • Comprehensive experiments on multiple benchmarks.
-Extensive experiments on the proposed RB-SCD and three public CD datasets demonstrate that MFDCD achieves superior performance and generalization ability in identifying complex semantic changes in traffic scenes.
-II Related Work
-II-A CD Datasets and Methods
-Existing datasets, as shown in Tab.
-I, such as  [6][7][8][28], are primarily designed for BCD tasks and cannot provide specific semantic change information.
-For SCD, datasets like  [9][10][11] have established some benchmarks.
-However, these datasets mainly focus on changes in land cover, with little attention paid to the changes in roads and bridges in traffic scenes.
+  • 1.
+    We introduce RB-SCD, a challenging benchmark dataset specifically designed for SCD of roads and bridges.
+It addresses the critical scarcity of domain-specific data by providing 11 fine-grained semantic change categories, enabling a more nuanced and practical analysis of traffic infrastructure evolution.
+  • 2.
+    We propose MFDCD, a framework that for the first time integrates multimodal (visual-textual) features in the frequency domain for CD.
+The framework overcomes conventional limitations through two key components: the DFC leverages wavelet transform to robustly model linear and heterogeneous semantic changes, and the TFF acts as a semantic reasoning engine to resolve high semantic ambiguity by aligning textual concepts with visual patterns in the frequency domain.
+  • 3.
+    Extensive experiments show the state-of-the-art performance of our MFDCD.
+Specifically, on our proposed RB-SCD dataset, it excels at identifying complex semantic changes in traffic scenes.
+On three public benchmarks for BCD, it shows superior generalization ability and achieves highly competitive results.
+2 Related Work
+2.1 CD Datasets and Methods
+Existing BCD datasets such as LEVIR [3], WHU [4], CDD [5], SYSU [6] are primarily designed for BCD tasks and cannot provide specific semantic change information.
+For SCD, datasets like Landsat [7], HRSCD [8], SECOND [9] mainly focus on changes in land cover, with little attention paid to the changes in roads and bridges in traffic scenes.
 In contrast, our RB-SCD dataset is a high-resolution, fine-grained SCD dataset specifically curated and annotated for road and bridge changes in traffic scenes.
-It captures a wide range of semantic change types and includes diverse geographical contexts, enabling precise and meaningful evaluation of CD algorithms in real-world urban transportation scenarios.
-By addressing both the lack of semantic detail and the absence of domain specificity in existing datasets, RB-SCD provides a valuable and targeted benchmark for infrastructure monitoring, intelligent transportation analysis, and traffic scene understanding.
-Figure 3: (a) Statistical distribution of pixels for semantic change categories. (b) Train set. (c) Test set.
-TABLE II: Specific subcategories in RB-SCD dataset.
-Categories Description
-Road street, highway, dirt road, walkway, railway, etc
-Bridge
-beam bridge, arch bridge, cable-stayed bridge,
-suspension bridge, etc.
-Bare land
-wasteland, sand patches area, unpaved open
-grounds, road or railway construction zones, etc.
-Grass
-urban lawn, green belt, wetland grass, riparian
-grass zone, meadow etc.
-Water
-ocean, lake, river, reservoir, pond, etc.
-Farmland agricultural land (with and without crops)
-Vegetation forest, shrub, urban vegetation buffer, etc.
-Building
-residential building, commercial building,
-industrial building, under-construction building
-(with clear structure), etc.
+It enables precise and meaningful evaluation of CD algorithms in traffic scenarios.
 In recent years, deep learning has significantly improved the accuracy of CD through its ability to extract relevant features from raw data.
-Changemask [32] decouples SCD into temporal semantic segmentation and BCD, and combines these two tasks into a unified encoder-transformer-decoder framework.
-MDENet [33] uses MDCM to capture object features in frequency and spatial domains across times, and MFGatt to reduce style impact, focusing on object location.
-JFRNet [34] introduces a three-branch network that takes joint bi-temporal features as unified input, enabling each branch to perceive the semantic relevance of the two time points without any additional interaction operations.
-DecoderFusion [35] enhances CD decoding features by leveraging the feature differences between two semantic segmentation decoding branches.
+C-3PO [21] reframes CD as a semantic segmentation task, proposing a new paradigm that adapts existing powerful segmentation networks for CD by introducing a dedicated module to learn different change types.
+S²PNet [22] enhances CD in complex VHR scenes by introducing a self-structured feature pyramid for better multi-scale object cognition.
+MHF²-Net [23] addresses the multi-scale object detection challenge in VHR CD by proposing a hierarchical feature fusion network that explicitly enhances high-frequency details.
+AGFormer [24] introduces a novel transformer-based model that mitigates classifier bias from class imbalance in CD by using an anchor-guided regularization strategy on a hypersphere to ensure better inter-class separability and intra-class balance.
 CdSC [25] introduces a cross-differential semantic consistency network that explores the differences in bi-temporal features while maintaining their semantic consistency.
@@
 However, these methods only consider single-modal data, i.e., images, while neglecting the rich semantic information embedded in multimodal data, thus encountering a performance bottleneck.
-II-B Multimodal Learning
+2.2 Multimodal Learning for Change Detection
 Recent studies have shown that multimodal information is beneficial for visual image analysis.
-LAVT [21] and VLT [38] integrate textual features into visual features through language encoders.
-Zhang et al.
-[22] propose a convolutional neural network method with an attention mechanism to enhance feature extraction for lidar data.
-Their approach leverages the elevation information from lidar data and the spatial-spectral information of hyperspectral data for multimodal feature fusion.
-MedCLIP [39] uses text prompts to generate segmentation for clinical scans, adopting a novel decoupled hard-negative noise contrastive estimation loss to fine-tune the BiomedCLIP model.
 ChangeCLIP [20] designs a novel CD framework that utilizes semantic information from image-text pairs and introduce an innovative differential feature compensation module to capture fine-grained semantic changes between them.
-II-C Frequency Transform in Deep Learning
+HFA-PANet [27] addresses the domain shift challenge in multimodal CD by proposing a non-siamese network that hierarchically aligns features from heterogeneous images to effectively capture their differences.
+AEKAN [28] introduces a novel unsupervised approach for multimodal CD by using a Siamese Kolmogorov-Arnold Network (KAN) AutoEncoder to effectively learn the latent commonality features between heterogeneous images.
+CFRL [29] presents an unsupervised framework for multimodal CD that learns comparable commonality features by cyclically reconstructing, re-encoding, and aligning features from heterogeneous images.
+While these pioneering works highlight the growing importance of multimodality, they primarily fall into two categories: visual-textual fusion in the spatial domain or feature alignment for heterogeneous image inputs.
+Our work carves out a distinct niche by being the first to introduce a visual-textual paradigm that operates within the frequency domain, specifically designed to resolve the deep semantic ambiguities found in complex, heterogeneous RS scenes
+2.3 Frequency-Domain Analysis in Deep Learning
 In image processing, the combination of CNN and frequency domain information has been widely applied in tasks such as image restoration, super-resolution, and semantic segmentation.
@@
 ARFFT [32] proposes a spatial-frequency fusion block to enhance the representation capability of Transformers and extend the receptive field across the entire image.
-XNet [43] utilizes wavelet transforms to transfer visual information to the frequency domain, leveraging high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks.
+XNet [33] utilizes high- and low-frequency information to design a framework supporting both fully supervised and semi-supervised tasks.
+Figure 3: Annotation examples on RB-SCD dataset. $T_{1}$ and $T_{2}$ represent the pre- and post-change images, respectively. The Label row shows the ground truth.
 Given the characteristics of RS images, such as high resolution and susceptibility to spectral variations caused by seasonal changes, using only frequency domain features often results in the loss of critical spatial information.
-Therefore, in this paper, we propose a novel architecture that models multimodal features in the frequency domain, effectively balancing the advantages of visual features, textual features, and frequency domain features, achieving promising results.
-Figure 4: Examples of dataset distribution spanning multiple representative geographical areas.
-Figure 5: Examples of road and bridge types, where the first two columns are roads, and the last three columns are bridges.
-III RB-SCD Dataset
-III-A Data Annotation and Analysis
-Rich Semantic Annotations: As shown in Fig.
+Therefore, in this paper, we model multimodal features in the frequency domain, effectively balancing the advantages of visual features, textual features, and frequency domain features, achieving promising results.
+3 RB-SCD Dataset
+As established, existing benchmarks are not suited for the task of road and bridge CD in traffic scenes.
+Therefore, our primary goal in creating the RB-SCD dataset is to provide detailed, fine-grained annotations that explicitly capture the semantic evolution of linear networks, while also covering a wide range of complex real-world scenes.
+3.1 Data Annotation and Analysis
+Rich Semantic Annotations.
+As shown in Fig.
 3 and Fig.
-3, unlike traditional BCD datasets, we have defined eight major categories (with specific subcategories detailed in Tab.
-II), including road, bridge, bare land, building, farmland, water, grass and vegetation.
-The RB-SCD dataset provides detailed semantic change annotations for 11 distinct types of semantic changes and a background category across these eight categories.
+4, unlike traditional BCD datasets, RB-SCD provides fine-grained annotations for 11 distinct semantic change types.
+These change annotations are constructed from eight fundamental land-cover classes, which are detailed in Tab.
+2.
+A critical contribution of RB-SCD is its fine-grained semantic granularity, which is not merely a detail but an essential feature for meaningful real-world applications.
+This granularity allows for a precise understanding of urban evolution.
+For instance, categories like farmland->road and vegetation->road directly track the impact of urbanization on
+Figure 4: Statistical analysis of the RB-SCD dataset.
+(a) Per-class pixel distribution for semantic change categories.
+(b-c) Semantic transition matrices for the training and test sets, respectively, showing the pixel counts for each from -> to change type.
+Here, bl, wat, veg, bd, fl, bri, and gra denote abbreviations for bare land, water, vegetation, building, farmland, bridge, and grass, respectively.
+Table 2: Specific subcategories of land-cover classes in RB-SCD dataset.
+Categories Description
+Road street, highway, dirt road, walkway, railway, etc.
+Bridge
+beam bridge, arch bridge, cable-stayed bridge, suspension bridge, etc.
+Bare land
+wasteland, sand patches area, unpaved open grounds, road or railway construction zones, etc.
+Grass
+urban lawn, green belt, wetland grass, riparian grass zone, meadow etc.
+Water
+ocean, lake, river, reservoir, pond, etc.
+Farmland agricultural land (with and without crops).
+Vegetation forest, shrub, urban vegetation buffer, etc.
+Building
+residential building, commercial building, industrial building, under-construction building
+(with clear structure), etc.
+Figure 5: The diverse geographic distribution of the RB-SCD dataset.
+The samples span multiple countries and continents, including varied urban and rural landscapes from China, Japan, America, Italy, Singapore, Australia, and Indonesia.
+This diversity ensures the dataset’s robustness and challenges models to generalize across different environmental conditions and infrastructure styles.
+Figure 6: Examples of road and bridge types, where the first four columns are roads, and the last three columns are bridges.
+ecological and agricultural land, providing critical data for environmental assessment.
+Conversely, changes such as road->grass or the rare but critical bridge->water capture infrastructure degradation or demolition, which is vital for safety monitoring and disaster response.
+Distinguishing between road widening (an implicit category captured through pixel-level changes) and building->road is crucial for urban renewal analysis and traffic management.
+All annotations are performed by a professional team in RS and underwent rigorous quality control to ensure accuracy and semantic consistency.
+It is worth noting that some changes, such as bridges directly turning into roads, are rarely observed in reality.
+Additionally, other types of changes, such as those between bridges and some other categories, occur in only a few samples, and we set their pixel number to 0 in the training process.
+Diverse Geographic Coverage.
+The image pairs are sourced from Google Earth, with a temporal coverage spanning from 2007 to 2022.
+As shown in Fig.
+5, following the principles in [34], we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set.
+The dataset covers various regions worldwide, including Chinese cities such as Hefei, Xi’an, Chongqing and Guangzhou, as well as representative areas from Japan, Singapore, Nigeria, Italy, America, and Indonesia, etc.
+This extensive geographic coverage ensures that the dataset encompasses traffic scenes under varying climatic conditions, geological structures, and urban planning contexts, thereby guaranteeing its diversity and representativeness.
+Multiple Road and Bridge Types.
+As shown in Fig.
+6, the dataset covers diverse road and bridge types across varied regions and urban contexts.
+Road categories range from urban main roads, secondary streets, and highways to elevated roads designed for congestion relief, while bridge types include large cable-stayed bridges, girder bridges, and pedestrian bridges.
+This diversity captures different engineering styles and traffic environments, providing robust support for training CD models across multiple scales and diverse scenarios.
+3.2 Dataset Limitations
+As shown in Fig.
+4, RB-SCD inevitably exhibits class imbalance across semantic change types.
+Common transitions (e.g., road widening, bridge construction) are well represented, while rarer cases (e.g., road repair, demolition) appear less frequently.
+We have used a weighted cross-entropy loss to mitigate the class imbalance issue in the RB-SCD dataset, which assigns higher weights to the losses of under-represented classes.
+We will further collect more samples of different types of changes in future work and employ techniques such as data augmentation and sample weighting to alleviate the model’s difficulty in detecting rare changes.
 Figure 7: Overview of our proposed Multimodal Frequency-Driven Change Detector (MFDCD).
-Firstly, the Dynamic Frequency Coupler (DFC) couples hierarchical visual features extracted by ResNet50 with high- and low-frequency components obtained via wavelet transform to enhance cross-temporal structural change perception.
-Meanwhile, the Textual Frequency Filter (TFF) transforms textual features extracted by a pre-trained CLIP model into the frequency domain using Fourier transform, applies graph-based filtering to extract salient components, and integrates them with visual features for collaborative representation learning.
-Finally, encoder features are fed into the decoder to obtain the final semantic change detection results.
-Changes such as road widening, new road construction, and bridge construction or modification are reflected through the corresponding semantic change types, supporting in-depth exploration of the intrinsic mechanisms of traffic scenario evolution.
-It is worth noting that some changes, such as bridges directly turning into roads, are rarely observed in reality.
-Additionally, other types of changes, such as those between bridges and some other categories, occur in only a few samples, and we set their pixel number to 0.
-All annotations are carried out by a professional team in RS and transportation, and underwent multiple rounds of rigorous review and quality control to ensure annotation accuracy and semantic consistency.
-Diverse Geographic Coverage: As shown in Fig.
-4, following the principles in [44], we collect 260 pairs of RS images with resolutions ranging from 270 × 534 to 7,215 × 4,366 (0.59 m/pixel), which can be cropped into 5,016 pairs of 256 × 256 patches, including 4,015 pairs for the training set and 1,001 pairs for the testing set.
-The dataset integrates satellite images obtained primarily from Google Earth, covering various regions worldwide, including Chinese cities such as Hefei, Xi’an, Chongqing and Guangzhou, as well as representative areas from Japan, Singapore, Nigeria, Italy, America, and Indonesia, etc.
-This extensive geographic coverage ensures that the dataset encompasses traffic scenes under varying climatic conditions, geological structures, and urban planning contexts, thereby guaranteeing its diversity and representativeness.
-Multiple Road and Bridge Types: As shown in Fig.
-5, the dataset encompasses various types of roads and bridges from different geographical regions and urban environments.
-The road categories include concrete and asphalt roads used for urban main roads, secondary roads, residential streets, and highways, comprehensively reflecting the common transportation routes found in cities.
-Additionally, the dataset contains elevated roads designed to alleviate congestion and improve traffic efficiency, which are typically found in densely populated or high-traffic urban areas.
-In terms of bridge structures, the dataset includes large cable-stayed bridges that span wide rivers or deep valleys, as well as girder bridges and pedestrian bridges used for short- to medium-span connections and pedestrian access.
-The dataset also includes other types of specialized roads and bridges, demonstrating a variety of engineering styles and applications.
-This richness in road and bridge types enables the dataset to represent complex traffic environments and provides robust support for training CD models across multiple scales and diverse scenarios.
-III-B Label Imbalance Analysis and Challenges
+The framework takes bi-temporal images ($T_{1},T_{2}$) as input.
+The Dynamic Frequency Coupler (DFC) processes visual features from a ResNet50 [35] backbone to robustly model linear and heterogeneous transitions.
+It employs an ASFF block to precisely delineate change boundaries and a BTFF block to model the overall change evolution, preventing fragmentation.
+Concurrently, the Textual Frequency Filter (TFF) processes CLIP-inferred textual descriptions to resolve high semantic ambiguity.
+It transforms text into the frequency domain and uses a graph filter to reason about the most relevant semantic concepts.
+Finally, the enhanced visual features from the DFC, guided by the semantic features from the TFF, are fed into a decoder to generate the final semantic change map.
+4 Methodology
+4.1 Overview
 As shown in Fig.
-3, our RB-SCD dataset exhibits a certain degree of sample imbalance In terms of semantic changes, there is an imbalance in the types of changes in the dataset, especially for fine-grained CD tasks.
-For some types of changes (e.g., road widening, bridge construction), there are more samples, while for other types of changes (e.g., road repair, demolition), there are fewer samples.
-This imbalance may lead to the model overfitting to the more common change types, while neglecting the rarer ones.
-Additionally, the imbalance between change and no-change samples is another issue, particularly for CD tasks.
-The no-change samples are relatively more abundant, which may lead to a decrease in the model’s sensitivity when processing change samples.
-To improve the accuracy and robustness of the model, we will further collect more samples of different types of changes in future work and employ techniques such as data augmentation and sample weighting to alleviate the model’s difficulty in detecting rare changes.
-IV Methodology
-IV-A Overview
+7, our MFDCD follows an encoder-decoder architecture.
+Given the bi-temporal images $T_{1}$ and $T_{2}$, we first employ a ResNet50 [35] backbone to extract multi-scale visual features $\mathbf{F}_{i}^{t}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2\right\})$.
+The core innovation consists of two parallel branches: a Dynamic Frequency Coupler (DFC) for linear and heterogeneous semantic change modeling and a Textual Frequency Filter (TFF) for semantic disambiguation.
+Finally, the enhanced features from these two branches are fused and passed to the decoder to generate the final semantic change map.
+4.2 Dynamic Frequency Coupler (DFC)
+To address the challenge of linear and heterogeneous semantic change transitions, we first employ the DWT operation.
+Unlike standard convolutions, as shown in Fig.
+2(a), DWT naturally decomposes features into two components that are crucial for this task: high-frequency components, which are sensitive to the detailed boundaries where heterogeneous land-cover transitions occur, and low-frequency components, which are ideal for capturing the long-range continuity and global structure of linear targets like roads.
+However, merely separating these frequency bands is insufficient; they must be intelligently integrated back into the feature hierarchy.
+Therefore, as shown in Fig.
+7, DFC incorporates two specialized fusion sub-modules: the Adaptive Sparse Frequency Fusion (ASFF) and the Bidirectional Temporal Frequency Fusion (BTFF) block.
+Adaptive Sparse Frequency Fusion (ASFF).
+ASFF block is designed for precise boundary delineation.
+It achieves this by fusing the distilled high-frequency details with the rich, fine-grained low-level spatial features from the backbone.
+Low-level features are rich in detailed information, while high-frequency features contain information about boundaries and abrupt changes, which naturally correspond to low-level features.
+Let $\mathbf{F}_{i}^{t}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}(i\in\left\{1,2\right\},t\in\left\{1,2\right\})$ be the low-level visual features from layer $i$ at time $t$.
+We first apply the DWT to obtain the frequency sub-bands:
+$\mathbf{LL}_{i}^{t},\mathbf{HL}_{i}^{t},\mathbf{LH}_{i}^{t},\mathbf{HH}_{i}^{t}=\mathcal{W}(\mathbf{F}_{i}^{t}),$ (1)
+where $\mathbf{LL}_{i}^{t}$ is the low-frequency band, $\mathbf{HL}_{i}^{t},\mathbf{LH}_{i}^{t},$ and $\mathbf{HH}_{i}^{t}$ signifying the high-frequency bands.
+$\mathcal{W}(\cdot)$ is the DWT operation.
+The high-frequency sub-bands are then fused into a unified high-frequency representation:
+$\mathbf{H}_{i}^{t}=C_{fuse}(Cat(\mathbf{HL}_{i}^{t},\mathbf{LH}_{i}^{t},\mathbf{HH}_{i}^{t})),$ (2)
+where $C_{fuse}(\cdot)$ is a convolutional block and $Cat(\cdot)$ is the concatenation operation.
+However, high-frequency features are highly susceptible to noise.
+Especially in the early stages of training, when the model is not yet well-optimized, these high-frequency features often contain significant clutter and irrelevant details.
+Therefore, we design a principled, sparsity-aware gating mechanism.
+$S(\cdot)$ is defined as a dynamic sparsity constraint function, which adaptively thresholds the high-frequency features based on the training progress:
+$\tilde{\mathbf{H}}_{i}^{t}=S(\mathbf{H}_{i}^{t},\lambda(p),g)=\begin{dcases}\medspace\medspace\mathbf{H}_{i}^{t},&{if}\quad|\mathbf{H}_{i}^{t}|>g\cdot\lambda(p)\\ \medspace\medspace\medspace\mathbf{0},&\quad\quad{otherwise}\end{dcases},$ (3)
+where $\lambda(p)=\lambda_{0}e^{-\delta p}$ is the dynamic sparsity strength, $\lambda_{0}$ is the initial sparsity strength, $\delta$ is the sparsity decay rate, $p$ is the current training iteration, and $g$ is a constant controlling the sparsity strength.
+Early in training, a strong constraint (large $\lambda(p)$) allows only the most salient boundary features to pass, ensuring stable learning.
+As training progresses, the constraint $\lambda(p)$ decays, gradually permitting finer details to be fused for precise boundary refinement.
+This coarse-to-fine approach is more robust than a fixed or learnable threshold, adapting to the model’s learning stage.
+Finally, we treat this refined high-frequency map as a spatially-aware gate to modulate the low-level features.
+The final fused feature representation $\mathbf{P}_{i}^{t}$ is generated by a fusion block $F_{\text{ASFF}}$:
+$\mathbf{P}_{i}^{t}=F_{\text{ASFF}}(\mathbf{F}_{i}^{t},\tilde{\mathbf{H}}_{i}^{t})=C_{out}(Cat(Up(\tilde{\mathbf{H}}_{i}^{t}),\mathbf{F}_{i}^{t})),$ (4)
+where $Up(\cdot)$ is a bilinear upsampling operator to match spatial dimensions, and $C_{out}$ is the output convolution layer.
+Bidirectional Temporal Frequency Fusion (BTFF).
+To robustly model the change evolution of linear structures and prevent fragmentation, the model must explicitly compare the structural information across two time points.
+The BTFF module is designed to achieve this via a cross-temporal structure injection mechanism.
+The core principle is to directly infuse the low-frequency structural representation from one time point into the high-level feature space of the other, thereby forcing the network to reason about semantic features in the context of the opposing temporal structure.
+Let $\mathbf{F}_{i}^{t}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}(i\in\left\{3,4\right\},t\in\left\{1,2\right\})$ be the high-level features and $\mathbf{LL}_{i}^{t}$ be the low-frequency structural components at time $t$ for layer $i$.
+To compute the temporally-informed feature representation $\mathbf{P}_{i}^{1}$ for time $t=1$, we inject the structural information from time $t=2$.
+The low-frequency component from the second time point, $\mathbf{LL}_{i}^{2}$, represents a structural prior.
+To prepare this prior for fusion, it must be projected into a higher-dimensional space compatible with the semantic features.
+This is achieved by a projection function $\Phi_{L}(\cdot)$, composed of a convolutional block followed by an upsampling operator $Up(\cdot)$:
+$\widehat{\mathbf{LL}}_{i}^{2}=\Phi_{L}(\mathbf{LL}_{i}^{2})=Up(C_{proj}(\mathbf{LL}_{i}^{2})),$ (5)
+where $C_{proj}(\cdot)$ maps the low-frequency features to an intermediate representation.
+Next, the semantic feature space of the first time point is augmented by the structural prior from the second.
+We define an augmentation operator $\Psi(\cdot,\cdot)$ which concatenates the two representations and applies a non-linear fusion transformation $C_{trans}(\cdot)$ to model their complex interactions:
+$\mathbf{P}_{i}^{1}=\Psi(\mathbf{F}_{i}^{1},\widehat{\mathbf{LL}}_{i}^{2})=C_{trans}(Cat(\mathbf{F}_{i}^{1},\widehat{\mathbf{LL}}_{i}^{2}))+\mathbf{F}_{i}^{1},$ (6)
+where $\mathbf{P}_{i}^{1}$ represents the learned structural difference and contextual information derived from the cross-temporal comparison.
+The process is performed symmetrically to compute $\mathbf{P}_{i}^{2}$ by injecting $\mathbf{LL}_{i}^{1}$ into $\mathbf{F}_{i}^{2}$, thus creating a bidirectional flow of structural information:
+$\mathbf{P}_{i}^{2}=\Psi(\mathbf{F}_{i}^{2},\widehat{\mathbf{LL}}_{i}^{1})=C_{trans}(Cat(\mathbf{F}_{i}^{2},\widehat{\mathbf{LL}}_{i}^{1}))+\mathbf{F}_{i}^{2}.$ (7)
+This mechanism ensures that the model’s understanding of each time point is conditioned on the structural reality of the other, making it highly effective at modeling change regions.
+4.3 Textual Frequency Filter (TFF)
+While recent vision-language models like ChangeCLIP [20] leverage textual priors, their fusion in the noisy spatial-domain still struggles to resolve deep semantic ambiguities.
+To overcome this, our Textual Frequency Filter (TFF) operates directly in the frequency domain to achieve a more robust semantic alignment.
+The core idea is to transform textual embeddings into a global signature that separates stable semantic concepts from high-frequency noise, thus providing clearer guidance to disambiguate challenging visual scenes.
+Our empirical validation in Fig.
+8 confirms this insight.
+We visualize the magnitude spectra of CLIP-generated text embeddings for representative change and no-change scenarios.
+In the no-change scene (Fig.
+8(b)), the $T_{1}$ and $T_{2}$ spectra are nearly identical, demonstrating the stability of the spectral representation.
+Conversely, in the change scenes (Fig.
+8(a) and  8(c)), the spectra diverge significantly, as highlighted by the large spectral difference (gray area).
+This demonstrates that semantic change is effectively encoded as a quantifiable signal in the frequency domain.
+Therefore, we employ the DFT as the foundational tool within our TFF.
+Figure 8: Visualization of text embedding spectra in the frequency domain by using the discrete Fourier transform.
+The blue dashed line represents the spectrum of the $T_{1}$ image description, while the red solid line is the $T_{2}$ description.
+The gray shaded area highlights the absolute difference between the two spectra.
 As shown in Fig.
-6, our MFDCD mainly includes two components: (1) DFC module effectively couples high-level and low-level visual features with high-frequency and low-frequency spectral features.
-(2) TFF module enhances the text information obtained through zero-shot inference, explores the representation of text features in the frequency domain, and promotes the interaction and fusion of multimodal data.
-The details are as follows.
-IV-B DFC Module
-Given the bi-temporal images $T_{1}$ and $T_{2}$, ResNet50 [45] is used as the backbone network to extract the corresponding multi-scale features $F_{i}^{t}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2)\right\}$.
-The features from the first two layers of the network are considered as low-level features, while the features from the last two layers are considered as high-level features.
-The low-frequency and high-frequency components of these features are then obtained through wavelet transform, which allows for the separation of coarse and fine-grained information in both spatial and frequency domains.
-Adaptive Sparse Frequency Fusion: Low-level features are rich in detailed information, while high-frequency features contain information about edges, details, and abrupt changes, which naturally correspond to low-level features.
-However, high-frequency features are sensitive to noise and may amplify the influence of irrelevant regions.
-As shown in the upper-right corner of Fig.
-6, the ASFF block applies a novel sparsity constraint to the high-frequency features, which is dynamically adjusted during training iterations.
-By combining the sparsity-constrained high-frequency features with low-level features, both the image’s detailed information (such as edges and textures) and local structural information are preserved.
-This combination makes the network more responsive to CD, especially in areas with detailed changes.
-The process is as follows:
-$LL_{i}^{t},HL_{i}^{t},LH_{i}^{t},HH_{i}^{t}=\mathcal{W}(F_{i}^{t}),$ (1)
-where $t\in\left\{1,2\right\}$, $i\in\left\{1,2,3,4\right\}$, $\mathcal{W}(\cdot)$ represents the discrete wavelet transform.
-$LL_{i}^{t}$ is the low-frequency band, $HL_{i}^{t}$, $LH_{i}^{t}$, and $HH_{i}^{t}$ signifying the high-frequency bands.
-$H_{i}^{t}=Conv(Cat(HL_{i}^{t},LH_{i}^{t},HH_{i}^{t})),$ (2)
-$\widetilde{H_{i}^{t}}=\mathcal{S}(H_{i}^{t},\lambda(p),g)=\begin{Bmatrix}H_{i}^{t},&if\left|H_{i}^{t}\right|>g\cdot\lambda(p)\\ 0,&otherwise\end{Bmatrix},$ (3)
-$\lambda(p)=\lambda_{0}\cdot e^{-\gamma p},$ (4)
-$P_{i}^{t}=Conv(Cat(Up(\widetilde{H_{i}^{t}}),F_{i}^{t})),$ (5)
-where $t\in\left\{1,2\right\}$, $i\in\left\{1,2\right\}$.
-$\mathcal{S(\cdot)}$ denotes the sparsity constraint function, which is used to apply the sparsity operation, $\lambda(p)$ is the dynamic sparsity strength, $\lambda_{0}$ is the initial sparsity strength, $\gamma$ is the sparsity decay rate, $p$ is the current training iteration, and $g$ is a constant controlling the sparsity strength.
-$H_{i}^{t}$ is the original fused high-frequency feature and $\widetilde{H_{i}^{t}}$ is the high-frequency feature after applying the sparsity constraint.
-$P_{i}^{t}$ is the enhanced feature that combines high-frequency and low-level information.
-$Conv(\cdot)$ and $Up(\cdot)$ are convolution and bilinear upsampling operation, respectively.
-Bidirectional Temporal Frequency Fusion: As shown in the lower-right corner of Fig.
-6.
-Take the low-frequency features of time $T_{2}$ and the high-level features of time $T_{1}$ as an example, high-level features at time $T_{1}$ typically contain the semantic hierarchy of the image, enabling the expression of global structural changes in the image, such as road widening and bridge construction.
-Low-frequency features at time $T_{2}$ mainly focus on smooth regions or large-scale change information in the image.
-By combining high-level features from $T_{1}$ with low-frequency features from $T_{2}$, the network can capture global structural changes across time points while maintaining sensitivity to large-scale, smooth region changes.
-This models the bidirectional relationship of changes: local changes at $T_{1}$ may lead to semantic-level changes at $T_{2}$, while detailed changes at $T_{2}$ may arise from regional distribution adjustments at $T_{1}$.
-The details of BTFF block are as follows:
-$P_{i}^{1}=Conv(Cat[Up(LL_{i}^{2}),F_{i}^{1}]),$ (6)
-$P_{i}^{2}=Conv(Cat[Up(LL_{i}^{1}),F_{i}^{2}]),$ (7)
-where $i\in\left\{3,4\right\}$, $P_{i}^{t}$ is the enhanced feature that combines low-frequency information and high-level information.
-IV-C TFF Module
-As shown in Fig.
-6, we leverage the powerful zero-shot inference capability of the CLIP model to obtain textual information from RS images for multimodal modeling.
+7, we first leverage the powerful zero-shot inference capability of the CLIP [36] model to obtain textual information from RS images for multimodal modeling.
 For the categories inferred by the CLIP model, we select the top five categories with the highest probability values.
-For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ,” where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself.
-$E_{T}$ is a pre-trained CLIP text encoder that encodes the text embedding.
-Considering that discrete Fourier transform can help to reveal high-order semantic patterns in text, especially by identifying more complex semantic information through the descriptors used at different temporal points ($T_{1}$, $T_{2}$).
-Through Fourier transform, periodic or repetitive structures in the text can be extracted, which helps identify commonalities and differences in changes.
-We first use Fourier transform to convert the text embeddings from the text domain to the frequency domain:
-$\mathbf{T}_{freq}=\mathcal{F}(E_{T}(T_{d})),$ (8)
-where $\mathcal{F}(\cdot)$ is the discrete Fourier transform, $\mathbf{T}_{freq}\in\mathbb{C}^{B\times N\times C}$, $B$ represents the batch size, $N$ represents the number of frequency components, $C$ represents the dimension of each frequency component.
-We introduce a graph filter bank $K=[k_{1},k_{2},...k_{k}]$ for the textual features, which propagate information between the nodes of the graph to filter and refine the frequency information, thereby obtaining salient features related to the changes.
+For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ” where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself.
+We leverage a pre-trained CLIP text encoder [36], denoted as $E_{T}$, to obtain token-level embeddings for each textual description $T_{d}$.
+$\mathbf{E}^{t}=E_{T}(T_{d}).$ (8)
+This process yields a feature tensor $\mathbf{E}^{t}\in\mathbb{R}^{B\times N\times D}$, where $B$ is the batch size, $N$ is the sequence length (number of tokens), and $D$ is the feature dimension for each token.
+To transform this sequence of token features into the frequency domain, we apply the DFT along the sequence dimension $N$.
+This operator, denoted as $\mathcal{F}$, analyzes the semantic patterns across the entire sequence.
+For each sample $b$ in the batch and each feature dimension $d$, the $k$-th component of the resulting frequency spectrum $\mathbf{T}_{{freq}}\in\mathbb{C}^{B\times N\times D}$ is computed as:
+$\mathbf{T}_{{freq}}(b,k,d)=\mathcal{F}(\mathbf{E}^{t})(b,k,d)=\sum_{n=0}^{N-1}\mathbf{E}^{t}(b,n,d)\cdot e^{-j2\pi kn/N},$ (9)
+where $n$ is the index over the sequence of tokens, $n\in\{0,\dots,N-1\}$, $k$ is the index for the frequency components, $k\in\{0,\dots,N-1\}$, $d$ is the index for the feature dimensions, $d\in\{0,\dots,D-1\}$, $j$ is the imaginary unit.
+The resulting $\mathbf{T}_{{freq}}$ is a complex-valued tensor that encodes how semantic features are distributed across different frequencies of the text sequence.
+This allows us to capture global relational patterns and periodic structures, yielding a more robust representation for subsequent reasoning.
+However, these frequency components are not independent.
+Complex semantic concepts are often defined by global and potentially non-local relationships across the spectrum.
+To capture these dependencies, we model the frequency components as nodes and introduce a graph filter bank to act as a semantic reasoning engine.
+Specifically, we employ a fully connected graph, a design choice that imposes the weakest structural prior.
+This is crucial because key semantic concepts (e.g., bare land and road) often have strong, non-local relationships within our structured text descriptions, which would be overlooked by models assuming locality, such as sparse graphs based on token adjacency.
+This maximally flexible structure allows our graph filter bank to learn arbitrary combinations of frequencies that characterize specific change types, without being prematurely constrained by locality assumptions.
 For the frequency feature $\mathbf{T}_{freq}$, we split it into real and imaginary parts and concatenate them to construct node features.
-The feature matrix $\mathbf{X}\in\mathbb{R}^{B\times N\times 2C}$ is the collection of features for all nodes, which allows us to convert frequency domain information into node features for the graph.
-Assuming that each frequency component is connected to every other frequency component (fully connected graph), the edge set $\mathbf{E}\in\mathbb{Z}^{2\times\frac{N(N-1)}{2}}$ of the graph can be constructed by combining each pair of nodes:
-$\mathbf{X}=Cat[\mathfrak{R}(\mathbf{T}_{freq}),\mathfrak{I}(\mathbf{T}_{freq})],$ (9)
-$\mathbf{E}=\left\{(i,j)\mid 1\leq i\neq j\leq N\right\},$ (10)
-$\mathbf{E}^{(k+1)}=\sigma(\hat{\mathbf{A}}\mathbf{E}^{(k)}\mathbf{W}_{k}),$ (11)
-$\mathbf{E}^{f}=\sum_{k=1}^{k}\mathbf{E}^{(k)},$ (12)
-where $\mathfrak{R}(\mathbf{T}_{freq})$ and $\mathfrak{I}(\mathbf{T}_{freq})$ represent the real and imaginary parts of the frequency domain features, respectively.
-$\mathbf{E}^{(k)}$ represents the node features after the $k$-th filter, $\hat{\mathbf{A}}$ is the normalized adjacency matrix, which represents the relationships between nodes, $\mathbf{W}_{k}$ is the weight matrix of the graph convolution filter at the $k$-th layer, and $\sigma$ is the activation function, $\mathbf{E}^{f}$ is the weighted sum of the outputs from multiple filters.
-We use the text-image encoder $E_{VT}$ of the CLIP model to fuse $\mathbf{E}^{f}$ and the image features.
-After subtracting the results from the bi-temporal phases, the output is fed into the decoder [46] to obtain our CD results.
-IV-D Frequency Transform Function Analysis
-Why use discrete wavelet transform for image features and discrete Fourier transform for textual features?
-The advantage of wavelet transform lies in its ability to analyze images at different scales and locations, making it particularly suitable for multi-level, multi-scale image analysis tasks.
-As shown in Fig.
-2, in remote sensing image CD, wavelet transform helps to extract high-frequency features (e.g., detailed changes) and low-frequency features (e.g., overall structure), aiding in the capture of changes at different spatiotemporal scales in the image.
-Considering that text is typically discrete sequential data, its changes are often global or frequency-related.
-Therefore, on the other hand, the Fourier transform can effectively reveal periodic semantic patterns in the text, which contributes to identify patterns and trends in text changes.
-V Experiments and Results
-V-A Public CD Datasets
-In addition to using our proposed dataset, as shown in Tab.
-III, we have employed three public CD datasets to evaluate the robustness of our model.
-The LEVIR-CD [7] dataset consists of 637 pairs of RS images with a resolution of 1,024 × 1,024.
-It mainly focuses on building area changes, such as building growth.
-The WHU-CD [6] dataset comprises a pair of high-resolution RS images with a size of 32,507 × 15,354 pixels.
-It covers areas before and after an earthquake and primarily includes building renovations and additions.
-The SYSU-CD [8] dataset is a collection of 20,000 pairs of high-resolution aerial images, each with a size of 256 × 256 pixels.
-The images are captured in the Hong Kong area and include urban construction, suburban expansion, and vegetation changes, etc.
-TABLE III: Public CD datasets used in the experiments.
-Datasets Train/ Val/ Test Image type Space resolution
-LEVIR-CD 7120/ 1024/ 2048 Satellite images 0.5 m/pixel
-SYSU-CD 12000/ 4000/ 4000 Aerial images 0.5 m/pixel
-WHU-CD 5947/ 743/ 744 Aerial images 0.2 m/pixel
-TABLE IV: Comparison of quantitative performance for semantic change detection on RB-SCD dataset.
+The feature matrix $\mathbf{X}$ is the collection of features for all nodes, which allows us to convert frequency domain information into node features for the graph:
+$\mathbf{X}=Cat(e(\mathbf{T}_{freq}),Im(\mathbf{T}_{freq})),\mathbf{X}\in\mathbb{R}^{B\times N\times 2C},$ (10)
+where $Re(\mathbf{T}_{freq})$ and $Im(\mathbf{T}_{freq})$ represent the real and imaginary parts of the frequency domain features, respectively.
+The edge set $\mathbf{E}$ can be constructed by combining each pair of nodes:
+$\mathbf{E}=\left\{(i,j)\mid 1\leq i\neq j\leq N\right\},\mathbf{E}\in\mathbb{Z}^{2\times\frac{N(N-1)}{2}}.$ (11)
+The graph filter bank is denoted as $\mathcal{K}=\{k_{1},k_{2},\dots,k_{L}\}$ and implemented as a set of $L$ parallel layers, $\{\mathcal{G}_{l}\}_{l=1}^{L}$, where each layer $\mathcal{G}_{l}(\cdot)$ corresponds to a filter $k_{l}$.
+Unlike a 1D CNN that captures only local dependencies, a graph filter performs relational reasoning over the entire spectral structure.
+Its message-passing mechanism is designed to learn a filter function that enhances or suppresses global spectral patterns associated with specific semantic changes.
+Each filter then is parameterized by its own weight matrix $\mathbf{W}_{l}$ and learns to capture a specific type of semantic relationship from the node features $\mathbf{X}$, thereby obtaining salient features related to the changes:
+$\mathbf{X}^{\prime}_{l}=\mathcal{G}_{l}(\mathbf{X},\hat{\mathbf{A}})=\sigma(\hat{\mathbf{A}}\mathbf{X}\mathbf{W}_{l}),$ (12)
+where $\mathbf{X}^{\prime}_{l}$ represents the node features after the $l$-th filter, $\hat{\mathbf{A}}$ is the normalized adjacency matrix derived from $\mathbf{E}$ and $\sigma$ is the activation function.
+The outputs from all filters in the bank $\mathcal{K}$ are then aggregated to form a comprehensive, refined semantic representation $\mathbf{X}_{T}$:
+[…]
+Let $V_{ct}$ be the spatial visual context tokens from the image backbone.
+A context decoder ${E}_{\text{VT}}$, aligns the refined frequency-domain semantics $\mathbf{X}_{T}$ with this spatial visual evidence to compute a visually-informed feature:
+$\mathbf{X}_{T}^{refined}={E}_{\text{VT}}(V_{ct},\mathbf{X}_{T}).$ (14)
+By analyzing textual semantics in the frequency domain and aligning them with spatial visual features, TFF effectively resolves semantic ambiguities and provides reliable guidance for change prediction.
+5 Experiments and Results
+5.1 CD Datasets
+In addition to using our proposed dataset, we have employed three public CD datasets to evaluate the robustness of our model.
+The LEVIR-CD [3] dataset consists of 637 pairs of 1,024 × 1,024 satellite images with a spatial resolution of 0.5 m/pixel.
+It primarily focuses on building area changes, and is split into 7,120/1,024/2,048 pairs for training, validation, and testing.
+The SYSU-CD [6] dataset is a collection of 20,000 pairs of 256 × 256 high-resolution aerial images, also at 0.5 m/pixel resolution.
+Captured in the Hong Kong area, it includes diverse changes like urban construction and suburban expansion, and is divided into 12,000/4,000/4,000 pairs for train/val/test.
+The WHU-CD [4] dataset comprises a pair of large-format aerial images with a size of 32,507 × 15,354 and a higher resolution of 0.2 m/pixel.
+It covers areas before and after an earthquake, primarily including building renovations and additions.
+Following the standard split, it provides 5,947/743/744 image patches.
+5.2 Implementation Details
+Our experiments are conducted on a single GeForce RTX 3090 GPU.
+The AdamW is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01.
+The maximum number of iterations is set to 40,000.
+For data augmentation, RandomFlip and PhotoMetricDistortion are applied.
+The constant $g$ in the ASFF block is set to 0.1, the decay rate $\delta$ is set to 0.0001, and $\lambda_{0}$ is set to 1.
+We use a weighted cross-entropy loss to mitigate the class imbalance issue in the RB-SCD dataset, which assigns a weight of 1.5 to the losses of under-represented classes, while more common classes retain a weight of 1.0, thereby increasing the penalty for misclassifying rarer change types.
+To evaluate the performance of our experiments, we adopt four widely used metrics: recall (Rec), precision (Pre), F1-score (F1) and intersection over union (IoU), where IoU and F1 are the primary metrics, while the others are auxiliary metrics.
+Table 3: Comparison of quantitative performance for semantic change detection on RB-SCD dataset.
 All metrics are expressed as percentages (%).
@@
 building->road 36.56 55.57 43.44 69.78 36.10 53.05 41.28 74.18 39.83 56.97 51.63 63.54 40.46 57.61 44.69 81.04 45.44 62.49 53.0675.98
-farmland->road 63.36 73.55 77.80 77.34 63.80 77.90 78.78 86.61 65.38 79.07 75.87 82.55 65.02 78.80 81.30 76.45 68.40 81.24 79.21 83.38
+farmland->road 63.36 73.55 77.80 77.34 63.80 77.90 70.78 86.61 65.38 79.07 75.87 82.55 65.02 78.80 81.30 76.45 68.40 81.24 79.21 83.38
 grass->road 36.55 53.54 39.09 84.94 36.66 53.65 46.45 64.48 40.60 57.7548.26 71.89 36.55 53.54 39.09 84.94 42.49 59.6447.24 80.86
 road->grass 22.44 36.65 32.38 42.21 15.53 26.89 17.41 58.97 18.41 31.10 22.75 49.14 21.76 35.7524.2168.26 21.65 35.59 22.86 80.30
-bridge->water 46.51 63.49 49.92 87.19 79.36 82.60 74.25 93.06 83.40 90.9593.45 88.57 77.86 87.55 90.44 84.85 83.58 91.05 90.34 91.78
+bridge->water 46.51 63.49 49.92 87.19 70.36 82.60 74.25 93.06 83.40 90.9593.45 88.57 77.86 87.55 90.44 84.85 83.58 91.05 90.34 91.78
 mIoU 48.65 52.75 54.63 54.95 57.52
-TABLE V: Performance of the RS foundation model LWGANet on RB-SCD dataset. All metrics are expressed as percentages (%).
+5.3 Quantitative Assessment Results
+Results for Semantic Change Detection on RB-SCD Dataset.
+Tab.
+3 presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics.
+Across this diverse range of semantic changes, our method achieves significant advantages in both the IoU and F1 metrics.
+Specifically, for the easily detectable categories like water->bridge, bridge->water, water-> road and farmland->road, our model attains IoU scores of 86.83%, 83.58%, 71.68%, and
+Table 4: Performance of the RS foundation model LWGANet on RB-SCD dataset.
 Class LWGANet-L1 [38] LWGANet-L2 [38]
@@
 mIoU 55.12 55.71
-TABLE VI: Comparison of quantitative performance for binary change detection on RB-SCD dataset.
-All metrics are expressed as percentages (%).
-Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
+Table 5: Comparison of quantitative performance for binary change detection on RB-SCD.
 Model IoU F1 Rec Pre
-FC-diff [15] 57.81 73.26 69.63 77.30
-FC-ef [15] 59.05 74.26 69.05 80.31
-FC-cat [15] 60.34 75.26 71.99 78.85
 SNUNet [13] 62.90 77.22 72.47 82.64
-BIT [18] 62.21 76.71 70.97 83.69
+BIT [16] 62.21 76.71 70.97 83.47
 ChangeFormer [17] 61.88 76.45 69.69 84.66
@@
 Ours 65.83 79.39 74.13 85.47
-TABLE VII: Comparison of quantitative performance of various CD techniques on public datasets.
-All metrics are expressed as percentages (%).
-Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
+Table 6: Comparison of quantitative performance of various change detection techniques on public datasets.
 Results marked with * are obtained from the original papers.
-Model LEVIR-CD [7] SYSU-CD [8] WHU-CD [6]
+The ‘-’ symbol indicates that the corresponding data is not available in the source.
+Model LEVIR-CD SYSU-CD WHU-CD
 IoU  F1  Rec  Pre IoU  F1  Rec  Pre IoU  F1  Rec  Pre
-FC-diff [15] 75.18 85.83 84.59 87.12 42.05 59.20 46.72 80.80 69.48 81.99 78.90 89.14
-FC-ef [15] 71.80 83.59 82.27 84.94 55.83 71.65 65.41 79.22 58.53 71.65 65.41 79.22
-FC-cat [15] 76.98 86.99 85.64 88.38 60.68 75.53 71.38 80.21 72.24 83.87 84.22 83.53
 SNUNet [13] 82.20 90.23 88.97 91.54 66.73 80.04 79.79 80.30 76.02 86.38 88.95 83.95
@@
 ChangeFormer [17] 81.93 90.07 87.38 92.92 60.46 75.36 71.06 80.22 83.88 91.23 90.76 91.71
-ICIF [17] 79.86 88.80 86.85 90.84 61.38 76.07 73.31 91.94 78.27 87.81 84.06 91.96
+ICIF [14] 79.86 88.80 86.85 90.84 61.38 76.07 73.31 79.05 78.27 87.81 84.06 91.96
 TFI-GR [39] 83.03 90.73 89.53 91.95 71.38 83.30 81.16 85.56 88.55 93.92 92.65 95.24
@@
 USSFC-Net [15] 82.54 90.44 91.82 89.09 65.75 79.34 81.52 77.27 83.92 91.26 93.20 89.40
-M-Swin * [51] 83.57 91.05 90.05 92.07 -   -   -   - 86.09 92.94 91.37 94.56
 ELGC-Net [19] 83.52 91.02 89.78 92.30 65.34 79.03 77.26 80.89 75.56 86.08 82.15 90.40
 MDENet * [40] 83.63 91.0992.68 89.55 -   -   -   - 87.74 93.47 95.49 91.54
+MHF²Net * [23] 82.64 90.50    -      - -   -   -   - 87.14 93.13    -      -
 Ours 84.69 91.7192.31 91.11 72.63 84.15 85.60 82.74 90.43 94.97 96.44 93.55
-V-B Implementation Details
-Our experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU.
-The AdamW [52] is employed, with a learning rate of 0.0001 and a weight decay parameter of 0.01.
-Cross entropy loss is used as our loss function, and the maximum number of iterations is set to 40,000.
-For data augmentation, RandomFlip and PhotoMetricDistortion are applied.
-The constant $g$ in the ASFF block is set to 0.1, the decay rate $\gamma$ is set to 0.0001, and $\lambda_{0}$ is set to 1.
-To evaluate the performance of our experiments, we adopt four widely used metrics: recall (Rec), precision (Pre), F1-score (F1) and intersection over union (IoU), where IoU and F1 are the primary metrics, while the others are auxiliary metrics.
-The following are formulas for defining the evaluation metrics as mentioned above.
-$Rec=TP/(TP+FN),$ (13)
-$Pre=TP/(TP+FP),$ (14)
-$F1=2/(Rec^{-1}+Pre^{-1}),$ (15)
-$IoU=TP/(TP+FP+FN),$ (16)
-where TP, TN, FP, and FN denote the counts of true positives, true negatives, false positives, and false negatives, respectively.
-V-C Quantitative Assessment Results
-Results for Semantic Change Detection on RB-SCD Dataset: Tab.
-IV presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics.
-As shown in the table, we have defined 11 types of semantic changes along with one background category, aiming to comprehensively cover the diverse real-world scenarios of road and bridge changes.
-These change types include both relatively straightforward categories—such as water to bridge and farmland to road, which typically involve simple backgrounds or abundant training samples—and more challenging categories, such as road to vegetation and grass to road, which are characterized by complex backgrounds, ambiguous boundaries, and limited sample sizes.
-Across this diverse range of semantic changes, our method achieves significant advantages in both the IoU and F1 metrics.
-Specifically, for the easily detectable categories—water to bridge, bridge to water, water to road, and farmland to road—our model attains IoU scores of 86.83%, 83.58%, 71.68%, and 68.40%, respectively, significantly outperforming other comparison methods.
+68.40%, respectively, significantly outperforming other comparison methods.
 These results highlight the strong capability of our model in semantic discrimination and spatial reasoning.
-More importantly, our method also performs excellently on several challenging categories, including road to vegetation, road to grass, road to bare land, and grass to road, which suffer from sample scarcity and semantic ambiguity.
+More importantly, our method also performs excellently on several challenging categories, including road->vegetation, road->grass, road->bare land, and grass->road, which suffer from sample scarcity and semantic ambiguity.
 These types of changes are often affected by similar surface textures, seasonal variability, and complex contextual information.
@@
 Furthermore, we evaluate the performance of the latest RS foundation model, LWGANet [38], on our proposed dataset (as shown in Tab.
-V).
+4).
 The results further underscore the challenging nature of our dataset.
-By incorporating fine-grained and practically meaningful semantic change categories in traffic scenarios, our dataset provides a more rigorous and instructive benchmark for training and evaluating CD models in real-world applications.
-Results for Binary Change Detection on RB-SCD Dataset: As shown in Tab.
-VI, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics.
+Results for Binary Change Detection on RB-SCD Dataset.
+As shown in Tab.
+5, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics.
 In terms of the IoU metric, our model ranks first with a score of 65.83%, slightly outperforming ELGC-Net [19], which achieves 65.53%.
@@
 Since the F1 is a harmonic mean of Pre and Rec, this result suggests that MFDCD achieves an excellent balance between detection accuracy and completeness.
-For recall, USSFC-Net [49] achieves the highest score of 75.07% among all models, while our model follows closely with a recall of 74.13%, ranking second.
-This demonstrates that USSFC-Net has a strong ability to detect all change areas, and our model also exhibits competitive performance in capturing changed regions.
-In terms of precision, our model outperforms others with a score of 85.47%, while ELGC-Net follows at 85.15%.
-This implies that our model has a lower false positive rate when identifying change areas.
-Overall, the results of the BCD task further indicate that existing methods still have room for improvement when applied to our RB-SCD dataset.
-Results on Public CD Datasets: To further prove the validity of our model, we also compare our proposed MFDCD with thirteen state-of-the-art CD techniques.
+Results on Public CD Datasets.
+To further prove the validity of our model, we also compare our proposed MFDCD with ten state-of-the-art CD models.
 To ensure a fair comparison, we reproduce these methods using publicly available code with default parameters.
-Tab.
-VII presents the quantitative evaluations of various methods on three public CD datasets, including IoU, F1, Rec, and Pre metrics.
 As observed in Tab.
-VII, our method achieves outstanding results across all three publicly available datasets.
+6, our method achieves outstanding results across all three publicly available datasets.
 Specifically, on the WHU-CD dataset, our approach outperforms the second-best method by 2.69% and 1.5% in IoU and F1, respectively.
@@
 These results further demonstrate the effectiveness and robustness of our method.
-Figure 7: Qualitative visualization for semantic change detection on RB-SCD dataset. (a) $T_{1}$ images. (b) $T_{2}$ images. (c) HGINet. (d) CdSC. (e) DEFO. (f) ChangeCLIP. (g) Ours. (h) Label.
+Figure 9: Qualitative visualization for semantic change detection on RB-SCD. (a) $T_{1}$ images. (b) $T_{2}$ images. (c) HGINet. (d) CdSC. (e) DEFO. (f) ChangeCLIP. (g) Ours. (h) Label.
+5.4 Qualitative Analysis.
+Qualitative Visualization for BCD Results.
+Fig.
+9 presents a qualitative comparison of MFDCD against state-of-the-art methods on challenging examples from the RB-SCD dataset.
+The visualizations highlight the tangible benefits of our multimodal, frequency-driven design.
+Superior Delineation of Linear Transitions: The first two rows demonstrate the effectiveness of our DFC in modeling linear and heterogeneous change transitions.
+In the first row, where a new road cuts through grass, other methods produce fragmented or incomplete predictions, especially under the road’s shadow.
+In contrast, our MFDCD, empowered by the DFC’s ability to model the long-range continuity of linear structures via its BTFF block, generates a much more complete and coherent road segment.
+Similarly, in the second row (bare land->road), the DFC’s ASFF block, which is sensitive to high-frequency boundary details, allows for a sharper and more precise delineation of the road edges compared to other methods.
+Robustness to Semantic Ambiguity: The third and fourth rows showcase the critical role of our TFF in resolving high semantic ambiguity.
+The third-row case (vegetation->road and grass->road) is a classic example of a scene where multiple land covers transition into a road.
+While other methods produce a confusing mix of class predictions, our TFF, acting as a semantic reasoning engine, provides a strong contextual prior.
+This guides the model to a confident and correct interpretation of the scene, resulting in a cleaner and more accurate segmentation.
+The fourth row further illustrates this, our method correctly identifies the subtle water->road change (light green) that other methods, including the language-guided ChangeCLIP, largely miss, demonstrating the effectiveness of our frequency-domain semantic alignment.
+Handling Pseudo-Changes and Complex Scenes: The final row (water->bridge) highlights the robustness of our model.
+The clean, sharp prediction of the new bridge demonstrates the DFC’s ability to extract the essential structural form while being less sensitive to noise like water texture.
+Qualitative Visualization for BCD Results.
+As shown in Fig.
+10, it can be observed that in areas where buildings are newly constructed or modified, most models can capture at least part of the change, though the precision of boundary detection varies.
+In regions where there is a transition from bare land to road, our MFDCD and some others show better performance in accurately delineating the change area.
+Some methods, however, have varying degrees of missed detections.
+For instance, methods such as BIT [16] and ChangeFormer [17] seem to have less accurate delineation compared to our method and ELGC-Net [19].
+Some models, like ICIF [14] and TFI-GR [39], can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives).
+Other models may miss parts of the change.
+In areas where the change from bare land to road is affected by shadows, most models struggle.
+Except for our MFDCD, which shows relatively better performance in distinguishing true changes from pseudo-changes caused by shadows, other models either misclassify the shadow areas as changes (false positives in red) or miss the actual changes (false negatives in blue).
 Figure 10: Qualitative visualization for binary change detection on RB-SCD dataset.
@@
 The rendered colors represent true positives (white), false positives (red), true negatives (black), and false negatives (blue).
+Table 7: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets.
+Method DFC TFF RB-SCD LEVIR-CD SYSU-CD WHU-CD
+(a) 50.45 81.65 68.98 86.65
+(b) 55.89 83.25 71.10 89.28
+(c) 55.23 82.97 70.89 88.50
+(d) 57.52 84.69 72.63 90.43
+Table 8: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets.
+Method RB-SCD LEVIR-CD SYSU-CD WHU-CD
+(a) - 55.23 82.97 70.89 88.50
+(b) $\mathbf{F}_{i}^{1}+\mathbf{H_{i}^{2}},i\in\left\{1,2,3,4\right\}$ 55.84 83.52 71.42 89.34
+(c) $\mathbf{F}_{i}^{1}+\mathbf{H_{i}^{2}},\mathbf{F}_{i}^{2}+\mathbf{H_{i}^{1}},i\in\left\{1,2\right\}$ 55.70 83.39 71.23 88.90
+(d) $\mathbf{F}_{i}^{t}+\mathbf{LL}_{i}^{t},i\in\left\{3,4\right\}$ 55.86 83.51 71.30 89.24
+(e) ASFF 56.42 84.02 72.06 89.98
+(f) BTFF 56.87 83.90 71.85 89.77
+(g) TFIM +TFF 55.98 83.86 71.63 88.97
+(h) ASFF+BTFF (DFC) 57.52 84.69 72.63 90.43
+Table 9: Effectiveness of the TFF module when trained on RB-SCD and public CD datasets.
+Method RB-SCD LEVIR-CD SYSU-CD WHU-CD
+(a) - 55.89 83.25 71.10 89.28
+(b) DFT+MLP 56.31 83.59 71.42 89.57
+(c) DFT+1D CNN 56.47 83.73 71.62 89.68
+(d) DFT+GAT 56.67 83.93 71.69 89.75
+(e) CLIP Embedding + Concat 56.38 83.54 71.85 89.84
+(f) DFT + Graph Filter (TFF) 57.52 84.69 72.63 90.43
+Table 10: Effectiveness of graph filters with different numbers $l$ in TFF.
+Method $l$ RB-SCD LEVIR-CD SYSU-CD WHU-CD
+(a) - 55.89 83.25 71.10 89.28
+(b) 3 56.74 83.84 71.98 89.87
+(c) 4 57.33 84.13 72.31 90.10
+(d) 5 57.52 84.69 72.63 90.43
+(e) 6 57.40 84.26 72.25 90.38
+Table 11: Effectiveness of MFDCD with different backbones when trained on RB-SCD and public CD datasets.
+Backbone Params (M) FLOPs (G) RB-SCD LEVIR-CD SYSU-CD WHU-CD
+ResNet18 13.75 11.32 55.53 83.18 71.35 89.04
+ResNet34 22.80 16.54 56.31 83.78 72.10 89.73
+ResNet50 103.78 49.02 57.52 84.69 72.63 90.43
+5.5 Ablation Study
+To demonstrate the effectiveness of our proposed modules, we conduct extensive experiments on RB-SCD and three public datasets.
+Effectiveness of DFC and TFF Modules.
+As shown in Tab.
+7, adding the DFC and TFF to the baseline model increases the mIoU on the RB-SCD dataset by 5.44% and 4.78%, respectively.
+Additionally, the IoU on the LEVIR-CD, SYSU-CD, and WHU-CD datasets improved by 1.6%/2.12%/2.63% and 1.32%/1.91%/1.85%, respectively.
+Detailed Analysis on DFC Module.
+As shown in Tab.
+8, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module.
+Initially, we experiment with different coupling methods for ordinary features and frequency domain features.
+By comparing rows (b) and (c) with row (h), we can observe that ASFF and BTFF effectively dynamically couple visual features with frequency domain features, achieving the best results.
+Comparing rows (c) and (e), it can be seen that combining the low-level features at time $T_{1}$ with the corresponding high-frequency features at time $T_{2}$, due to semantic mismatches, results in worse performance compared to our designed ASFF.
+Comparing rows (d) and (f), it is clear that our designed BTFF captures the global structural changes across time, outperforming the coupling of high-level features and low-frequency features at the same time point.
+To further validate our design, we replace our DFC with TFIM [39] (row (g)), a strong module for spatial feature interaction.
+Although TFIM provides a notable improvement over the baseline, our frequency-domain DFC yields better performance, demonstrating the superiority of explicitly modeling structural features in the frequency domain for this task.
+Detailed Analysis on TFF Module.
+As shown in Tab.
+9, the MLP (row (b)), which processes each frequency component independently, performs the worst, highlighting the necessity of modeling relational information.
+The 1D CNN (row (c) captures only local relationships and offers a slight improvement.
+Notably, while the GAT adaptively learns edge weights (row (d)), our proposed method (row (e)) with a simple graph filter over a fixed fully connected graph still achieves the best results.
+This indicates long-range dependency across the entire spectrum is more effective than attempting to dynamically prune or re-weight connections To validate our frequency-domain approach, we benchmark our TFF module against a powerful CLIP Embedding + Concat [14] mechanism (row (e)) that fuses semantics in the spatial/token domain.
+It shows that while this mechanism improves upon the vision-only baseline, our TFF-based model (row (f)) achieves superior performance.
+This suggests that extracting a robust, global semantic signature via frequency analysis is more effective for disambiguating complex RS scenes than performing direct spatial alignment with text tokens.
+As shown in
 Figure 11: Heatmaps of our proposed modules.
 (a) and (g) Input images.
-(b) and (h) Low-level features from the backbone network.
+(b) and (h) Low-level features from the backbone.
 (c) and (i) High-level features from the backbone network.
@@
 (f) and (l) Feature maps after undergoing the TFF module.
-V-D Quantitative Analysis.
-As shown in Fig.
-7, we compared the visualized results of the proposed MFDCD with other state-of-the-art methods on RB-SCD dataset.
-From the figure, it can be observed that our method outperforms others in identifying small semantic change regions.
-For instance, as shown in the third row, our method better distinguishes between changes from grass to road and from vegetation to road—two types of changes with similar characteristics.
-Regarding pseudo-changes caused by background composition, such as in the fifth row, where a boat passing under a bridge with its bow color similar to the bridge’s edge color, both DEFO [35] and ChangeCLIP [23] made incorrect predictions, while our method accurately detected the changes.
-Additionally, the proposed dataset presents significant challenges.
-For example, in the first row, all methods struggle to detect semantic changes under road shadows.
-As shown in the second row, farmland in autumn and winter appears visually similar to bare land, and in the fourth row, farmland in spring and summer shares similar colors with water, resulting in low inter-class differences.
-In such challenging scenes, other methods show many false detections, while our method performs slightly better but still leaves room for improvement.
-When comparing the BCD results of each method with the ground truth in Fig.
-8, it can be observed that for the unchanged areas (black regions representing true negatives), most models show relatively accurate results, with fewer false positives.
-In contrast, for the changed areas, the performance varies among models.
-Some models exhibit noticeable missed detections (blue regions representing false negatives), indicating they fail to identify all actual changes.
-Many models are capable of detecting building—related changes, which are common in the dataset.
-For example, in areas where buildings are newly constructed or modified, most models can capture at least part of the change, though the precision of boundary detection varies.
-In regions where there is a transition from bare land to road, our MFDCD and some others show better performance in accurately delineating the change area.
-Some methods, however, have varying degrees of missed detections.
-For instance, methods such as BIT [18] and ChangeFormer [19] seem to have less accurate delineation compared to our method and ELGC-Net [50].
-In areas where buildings are under construction in one image and completed in another, different models have different performances.
-Some models, like ICIF [17] and TFI-GR [48], can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives).
-Other models may miss parts of the change.
-In areas where the change from bare land to road is affected by shadows (resulting in pseudo-changes), most models struggle.
-Except for our MFDCD, which shows relatively better performance in distinguishing true changes from pseudo-changes caused by shadows, other models either misclassify the shadow areas as changes (false positives in red) or miss the actual changes (false negatives in blue).
-TABLE VIII: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets.
-All values are reported as percentages (%).
-Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
-Method DFC TFF RB-SCD LEVIR -CD SYSU -CD WHU -CD
-(a) × × 50.45 81.65 68.98 86.65
-(b) × 55.89 83.25 71.10 89.28
-(c) × 55.23 82.97 70.89 88.50
-(d) 57.52 84.69 72.63 90.43
-TABLE IX: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets.
-All values are reported in percentage (%).
-Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
-Method RB-SCD LEVIR -CD SYSU -CD WHU -CD
-(a) - 55.23 82.97 70.89 88.50
-(b) $F_{i}^{1}+H_{i}^{2},i\in\left\{1,2,3,4\right\}$ 55.84 83.52 71.42 89.34
-(c) $F_{i}^{t}+LL_{i}^{t},i\in\left\{1,2,3,4\right\}$ 56.10 83.67 71.68 89.45
-(d) $F_{i}^{1}+H_{i}^{2},F_{i}^{2}+H_{i}^{1},i\in\left\{1,2\right\}$ 55.70 83.39 71.23 88.90
-(e) $F_{i}^{t}+LL_{i}^{t},i\in\left\{3,4\right\}$ 55.86 83.51 71.30 89.24
-(f) ASFF 56.42 84.02 72.06 89.98
-(g) BTFF 56.87 83.90 71.85 89.77
-(h) ASFF+BTFF 57.52 84.69 72.63 90.43
-TABLE X: Effectiveness of graph filters with different numbers $k$ in TFF.
-All values are reported in percentage (%).
-Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
-Method $k$ RB-SCD LEVIR-CD SYSU-CD WHU-CD
-(a) - 55.89 83.25 71.10 89.28
-(b) 3 56.74 83.84 71.98 89.87
-(c) 4 57.13 84.13 72.31 90.10
-(d) 5 57.52 84.69 72.63 90.43
-(e) 6 57.40 84.26 72.25 90.38
-TABLE XI: Effectiveness of MFDCD with Different Backbones when Trained on Four CD Datasets.
-All Values are Reported in Percentage (%).
-Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
-Backbone Params (M) FLOPs (G) RB -SCD LEVIR -CD SYSU -CD WHU -CD
-ResNet18 13.75 11.32 55.53 83.18 71.35 89.04
-ResNet34 22.80 16.54 56.31 83.78 72.10 89.73
-ResNet50 103.78 49.02 57.52 84.69 72.63 90.43
-V-E Ablation Study
-To demonstrate the effectiveness of our proposed modules, we conduct extensive experiments on RB-SCD and three public datasets.
-As shown in Tab.
-VIII, when the DFC module and TFF module are added to the baseline model, the mIoU on RB-SCD dataset increased by 5.44% and 4.78%, respectively.
-Additionally, the IoU on the LEVIR-CD, SYSU-CD, and WHU-CD datasets improved by 1.6%/2.12%/2.63% and 1.32%/1.91%/1.85%, respectively.
-As shown in Tab.
-IX, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module.
-Initially, we experiment with different coupling methods for ordinary features and frequency domain features.
-By comparing rows (b) and (c) with row (h), we can observe that ASFF and BTFF effectively dynamically couple visual features with frequency domain features, achieving the best results.
-Comparing rows (d) and (f), it can be seen that combining the low-level features at time $T_{1}$ with the corresponding high-frequency features at time $T_{2}$, due to semantic mismatches, results in worse performance compared to our designed ASFF.
-Comparing rows (e) and (g), it is clear that our designed BTFF captures the global structural changes across time, outperforming the coupling of high-level features and low-frequency features at the same time point.
-As shown in Tab.
-X, we conduct ablation experiments on the number of layers in the frequency filter of TFF.
-As the value of $k$ increased, the model’s performance improved.
+Figure 12: Hyperparameter sensitivity analysis of the dynamic sparsity constraint in ASFF.
+We evaluate the impact of the initial sparsity $\lambda_{0}$ (left), the decay rate $\delta$ (middle) and the constant $g$ (right) on the mIoU performance on RB-SCD.
+Tab.
+10, we conduct ablation experiments on the number of layers in the frequency filter of TFF.
+As the value of $l$ increased, the model’s performance improved.
 However, when the number of filter layers reached six, the performance stabilized or slightly declined.
-Subsequent experiments are based on the optimal setting, $k=5$.
-To investigate the impact of backbone networks on the performance of the MFDD model in the CD tasks, as shown in Tab.
-XI, in terms of parameters and computational complexity, ResNet18 has the lowest cost (13.75M parameters, 11.32G FLOPs), while ResNet50 has the highest (103.78M parameters, 49.02G FLOPs).
+Subsequent experiments are based on the optimal setting, $l=5$.
+Backbone Analysis.
+To investigate the impact of backbone networks of the MFDCD model, as shown in Tab.
+11, in terms of parameters and computational complexity, ResNet18 has the lowest cost, while ResNet50 has the highest.
 In terms of detection accuracy, for all four datasets, the deeper the backbone network, the higher the accuracy.
@@
 In this paper, ResNet50 is selected as the backbone network.
+Visualization Analysis.
 To more clearly illustrate the influence of our proposed modules, we visualized heatmaps for these modules.
 As depicted in columns (b)/(h) and (c)/(i) of Fig.
-9, the low-level and high-level features from the backbone network typically exhibit certain characteristics.
+11, the low-level and high-level features from the backbone network typically exhibit certain characteristics.
 The low-level features may contain fine-grained but scattered information, while the high-level features, though more abstract, might lose some detailed spatial information.
 In columns (d)/(j), (e)/(k), and (f)/(l), after undergoing the ASFF block, BTTF block and TFF module respectively, the feature maps show progressive changes.
-The ASFF block helps to fuse features in a reasonable way, making the information more integrated.
-The BTTF block further refines the features, enhancing the relevant change information.
-The TFF module finally outputs feature maps that can better highlight the change regions, eliminating some irrelevant noise and enabling the acquisition of more distinct change-related information, which showcases the efficacy of the introduced modules in feature processing for CD tasks.
-VI Conclusion
-In this work, we introduce RB-SCD dataset, a SCD dataset specifically designed for complex traffic scenarios.
-The dataset features fine-grained pixel-level annotations across 11 semantic change categories and one background class, encompassing a wide range of road and bridge types.
-It also covers diverse geographic regions across multiple countries and urban environments, making it become a comprehensive and challenging benchmarks to date for road and bridge change analysis in RS imagery.
+The ASFF block helps to fuse features in a reasonable way, precisely delineate the boundaries.
+The BTTF block further maintains the continuity and integrity of linear structures.
+The TFF module finally outputs feature maps that can better highlight the change regions, eliminating some irrelevant noise and resolving semantic ambiguities.
+Hyperparameter Analysis.
+As shown in Fig.
+12, we conduct a sensitivity analysis on the key hyperparameters $\lambda_{0}$, $\delta$ and $g$ of our dynamic sparsity constraint.
+The performance remains stable across a reasonable range of values, demonstrating the robustness of our approach.
+Our chosen values ($\lambda_{0}$=1.0, $\delta$=0.0001,gg=0.1) provide the best performance.
+This dynamic mechanism proves more effective than a fixed threshold, as it adapts the level of detail refinement to the model’s learning stage, ensuring both stability in early training and precision in later stages.
+6 Conclusion
+In this work, we introduce RB-SCD, a SCD dataset specifically designed for complex traffic scenarios.
+RB-SCD provides comprehensive fine-grained pixel-level annotations across 11 semantic change categories, encompassing a wide range of road and bridge types.
+It also covers diverse geographic regions across multiple countries and urban environments, making it a challenging benchmark for road and bridge change analysis in traffic scenes.
 Notably, RB-SCD supports both BCD and SCD tasks, offering flexibility for different application scenarios.
-By incorporating both common and rare semantic transformations, RB-SCD provides a valuable resource for the development and evaluation of CD models in traffic contexts.
-To address the unique challenges presented by this dataset, we propose a novel multimodal CD framework, MFDCD, which leverages both spatial and frequency domain representations to enhance feature learning.
-By introducing frequency-domain analysis into the CD pipeline, MFDCD captures subtle semantic differences and structural changes that may be overlooked by conventional spatial-domain approaches.
-Our method also incorporates an effective fusion mechanism to integrate multimodal features, enabling more accurate detection of complex changes such as occluded road regions.
-Extensive experiments conducted on the RB-SCD dataset and public CD datasets demonstrate the superior performance of MFDCD across various metrics when compared to state-of-the-art CD methods.
-The results highlight the robustness, generalization ability, and semantic discrimination power of our approach in both straightforward and highly ambiguous change scenarios.
-In future work, we plan to extend our exploration of road and bridge change detection by integrating additional modalities such as elevation data or temporal sequences, and further investigating the role of frequency-domain representations in improving the interpretability and adaptability of CD models.
-As the label imbalance analysis and challenges for RB-SCD dataset discussed in Sec.
-III-B, we also plan to expand the scale of the dataset and enrich it with more detailed spatiotemporal annotations to support broader research tasks such as urban infrastructure evolution and intelligent transportation systems.
-Furthermore, while MFDCD achieves strong performance, it is not yet lightweight enough for deployment in resource-constrained environments.
-Future research will focus on model compression and efficient architecture design to improve its applicability in real-time and edge scenarios.
+We believe this benchmark represents an important first step towards advancing SCD of roads and bridges, and we envision future extensions to larger-scale and more balanced datasets.
+In addition, we propose the MFDCD framework, which pioneers a new paradigm by integrating multimodal features in the frequency domain.
+It tackles the core challenges with two specialized components: the DFC module, which leverages wavelet analysis to robustly model the evolution of linear structures, and the TFF module, which acts as a semantic reasoning engine to disambiguate visually similar land covers.
+Extensive experiments on RB-SCD and three public datasets validate the superiority of MFDCD over state-of-the-art methods, highlighting its robustness, generalization, and strong discrimination power.
+Limitation and Feature Work.
+Nevertheless, we acknowledge its limitations.
+The performance of MFDCD still degrades under extreme shadow or for rare, under-represented change classes, and the framework is not yet lightweight enough for deployment in resource-constrained scenarios.
+Future work will therefore focus on three directions: (1) expanding RB-SCD with more rare-class samples and temporal sequences; (2) extending our frequency-driven paradigm to topology-aware change reasoning for better connectivity analysis and pseudo-intersection recognition; and (3) developing lightweight model variants via compression and efficient architecture design to enhance applicability in real-time and edge-computing scenarios.
+Acknowledgement
+This work was supported in part by the NSFC Key Project of Joint Fund for Enterprise Innovation and Development under Grant U24A20342, and in part by the National Natural Science Foundation of China under Grant 62576006 and 61976004.